### v3

# Scaling Laws for Black-box Adversarial Attacks

Abstract
Adversarial examples usually exhibit good cross-model transferability, enabling attacks on black-box models with limited information about their architectures and parameters, which are highly threatening in commercial black-box scenarios. Model ensembling is an effective strategy to improve the transferability of adversarial examples by attacking multiple surrogate models. However, since prior studies usually adopt few models in the ensemble, there remains an open question of whether scaling the number of models can further improve black-box attacks. Inspired by the scaling law of large foundation models, we investigate the scaling laws of black-box adversarial attacks in this work. Through theoretical analysis and empirical evaluations, we conclude with clear scaling laws that using more surrogate models enhances adversarial transferability. Comprehensive experiments verify the claims on standard image classifiers, diverse defended models and multimodal large language models using various adversarial attack methods. Specifically, by scaling laws, we achieve 90%+ transfer attack success rate on even proprietary models like GPT-4o. Further visualization indicates that scaling laws exists on the interpretability and semantics of adversarial perturbations.
22footnotetext: Our code is available at https://github.com/liuchuan22/ScaleUpAttack.
1 Introduction
Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread applications [54]. However, deep neural networks (DNNs) remain susceptible to adversarial examples [35, 73] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions. Adversarial attacks have garnered considerable interest due to their implications for understanding DNN functionality [21], assessing model robustness [98], and informing the development of more secure algorithms [65].
Figure 1: Key observation of scaling laws for the transfer-based black-box attack. By scaling the number of surrogate models in the ensemble, we observe that the attack success rate and the discriminative loss on the target model follow a clear scaling law.
Adversarial attacks are typically categorized into white-box and black-box attacks based on whether the attacker has knowledge of the architecture and parameters of the target model. Given the greater risks with commercial services, e.g., GPT-4 [68], black-box attack methodologies [60, 22] have attracted growing attention. One prominent approach within black-box attacks is the transfer-based attack, which approximately optimizes the adversarial examples against the black-box model by minimizing the discriminative loss on a selection of predefined surrogate models. Researchers have devised advanced optimization algorithms [65, 81, 57] and data augmentation techniques [63, 23] in transfer-based attacks. Model ensembling [60, 22] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the transferability of adversarial examples. By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [22, 9].
Despite the progress in ensemble attack algorithms, the number of surrogate models used in previous research is often limited [9, 97, 94]. Recent advancement in large language models shows that scaling training data can significantly improve the generalization ability [48]. Since the transfer-based attack is also a generalization problem that expects adversarial examples from surrogate models to generalize to target models, while the training data are surrogate models rather than text or images [22, 9], this equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attacks by scaling the number of surrogate models, akin to the scaling of data and model parameters in large language models like GPT?
In this work, we investigate the scaling laws for black-box adversarial attacks. Theoretically, we provide an asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec. 3.1. Empirically, we show that the attack success rate and the loss are converging exponentially. We verify this by comprehensive experiments over traditional image classifiers (Sec. 3.2), defended models including adversarial training [65, 86] and DiffPure [67] (Sec. 3.3), and modern multimodal large language models (Sec. 3.4 and Sec. 4) using various attack methods (Sec. 5). These experimental results demonstrate the universality of scaling laws in transfer-based attacks.
More concretely, through scaling laws, we achieve over 30% improvement against vision classifiers and over 10% improvement against defended models. By extending the surrogate models to a group of CLIP [70] vision encoders, we achieve over 90% attack success rate against GPT-4o and GPT-4-Turbo, and over 80% against other commercial MLLMs. This highlights the vulnerabilities of these models to perturbations and the significance of scaling laws for ensemble attacks. We also demonstrate clear scaling laws on the interpretability and semantics of adversarial perturbations. Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than designing intricate algorithms for a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy MLLMs.
2 Backgrounds
In this section, we provide necessary backgrounds, including problem formulation and algorithms.
2.1 Problem Formulation
In this work, we focus on the challenging targeted attack setting. Let $\mathcal{F}$ denote the set of all image classifiers. Given a natural image $\bm{x}_{\text{nat}}$ and target class $\bm{y}$, the goal of targeted transfer-based attacks is to craft an adversarial example $\bm{x}$ that is misclassified to $\bm{y}$ by all models in $\mathcal{F}$. It can be formulated as a constrained optimization problem:
$\min_{\bm{x}}\mathbb{E}_{f\in\mathcal{F}}\left[\mathcal{L}(f(\bm{x}),\bm{y})\right],\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (1)
where $\mathcal{L}$ is the loss function (e.g., cross-entropy loss), and we consider adversarial perturbations in $\ell_{\infty}$ norm.
Generally, we take a limited set of “training” classifiers $\mathcal{F}_{T}:=\{f_{i}\}_{i=1}^{T}\subset\mathcal{F}$ as model ensemble to approximate problem (1) and therefore the optimization objective becomes:
$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (2)
where $T=\left\lvert\mathcal{F}_{T}\right\rvert$ is the cardinality of model ensemble.
2.2 Prerequisite Algorithms
Researchers have been developing sophisticated algorithms for ensemble attacks, and we select the state-of-the-art algorithms from them to conduct our experiments.
MI-FGSM [22] is a pioneering method of transfer-based attacks. It stabilizes the gradient update directions by integrating a momentum term into the iteration:
$\bm{g}_{t+1}=\mu\cdot\bm{g}_{t}+\frac{\nabla_{\bm{x}}\mathcal{L}(\bm{x_{t}},\bm{y})}{\left\lVert\nabla_{\bm{x}}\mathcal{L}(\bm{x_{t}},\bm{y})\right\rVert_{1}},$ (3)
and updates $\bm{x}$ by $\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1})$.
Common Weakness Method (CWA) [9] seeks to identify common vulnerabilities across surrogate models by encouraging flatter loss landscapes and closer proximity between the local optima of these models. It minimizes the second-order Taylor series:
$\frac{1}{T}\sum_{i=1}^{T}\left[\mathcal{L}(f_{i}(\bm{p}_{i},\bm{y})+\frac{1}{2}(\bm{x}-\bm{p}_{i})^{\top}H_{i}(\bm{x}-\bm{p}_{i})\right],$ (4)
where $\bm{p}_{i}$ denotes the closest optimum of model $f_{i}$ to $\bm{x}$ and $H_{i}$ is the Hessian matrix at $\bm{p}_{i}$.
Spectrum Simulation Attack (SSA) [63] applies a spectrum transformation $\mathcal{T}$ to the $\bm{x}$ and calculate gradients with respective to $\mathcal{T}(\bm{x})$ instead of $\bm{x}$, and can be combined with CWA to form SSA-CWA algorithm (pseudocode is provided in Appendix C). As shown in Table 1, SSA-CWA gives the best attack success rate. Therefore, in this paper, we use SSA-CWA as the core algorithm. We also explore the scaling laws of other algorithms in Sec. 5.1.
ASR Target: Car Target: Ship
Methods MI-CWA SSA-CWA MI-CWA SSA-CWA
Swin-v2 0.44 0.65 0.36 0.62
ConvNeXt 0.52 0.58 0.53 0.62
ResNeXt50 0.69 0.79 0.66 0.77
Table 1: Comparison of attack success rate (ASR) between SSA and MI algorithms. We follow the experiment settings in Sec. 3.2, and measure the ASR on 2 target classes and 3 black-box models.
Figure 2: Comparisons between CWA and MI-FGSM methods. We draw inputs $\bm{x}$ randomly from the dataset in Sec. 3.2 and measure the averaged gradient magnitudes on $\bm{x}$ after 10 steps of iterations. The results are recorded across $T$ from $1$ to $30$. On the right side, there are the visualizations of gradients of $T=30$.
Mitigating Gradient Conflict. We discover a new phenomenon of gradient conflict when ensembling a large number of surrogates. As shown in Fig. 2, as the number of models increases, the conflict between gradients of different models using common optimizers (e.g., MI-FGSM) becomes more severe, eventually reducing the ensemble gradient to zero and resulting in complete optimization stagnation. We conjecture this is because the gradients of normally trained models commonly lack semantics and have random directions; thus, naïve ensemble methods that directly average these gradients lead to strong gradient conflicts, significantly impacting optimization. To address this problem, we suggest to use optimizers that reduce the variance of gradients (e.g., SVRE [92]) or encourage the cosine similarity of gradients (e.g., CWA [9]). As demonstrated in Fig. 2, these methods fully avoid this issue. We analyze the gradient conflict on different attackers in more detail in Sec. 5.1.
3 Scaling Laws for Adversarial Attacks
In this section, we present our key findings on scaling laws for black-box adversarial attacks. Sec. 3.1 introduces fundamental concepts and theorems that give us motivations for the study of scaling laws. Sec. 3.2 contains our experiment settings and derivations of empirical scaling laws. Sec. 3.3 demonstrates clear scaling laws when attacking common defended models. Sec. 3.4 shows the results over multimodal large language models.
3.1 Motivations
First, we observe that in most practical scenarios, $\bm{x}_{\text{nat}}$ and $\bm{y}$ are predefined and fixed. Consequently, Eq. (1) and Eq. (2) can be considered as functions of variable $\bm{x}$ only.
Definition 1.
Define $\bm{x}^{*}$ and $\hat{\bm{x}}$ to be the minimizers of Eq. (1) and Eq. (2), respectively, which can be written as
$\bm{x}^{*}=\arg\min_{x}\mathbb{E}_{f\in\mathcal{F}}[\mathcal{L}(f(\bm{x}),\bm{y})],$ (5)
$\hat{\bm{x}}=\arg\min_{x}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}),$
with the same $\ell_{\infty}$ norm constraints given in Eq. (1) and  (2).
Definition 2.
Define $\mathbb{L}(\cdot)$ as the population loss in Eq. (1), and $\mathbb{L}(f_{i}(\cdot))$ as the loss on one surrogate model $f_{i}$:
$\mathbb{L}(\bm{x})=\mathbb{E}_{f\in\mathcal{F}}[\mathcal{L}(f(\bm{x}),\bm{y})],$ (6)
$\mathbb{L}(f_{i}(\bm{x}))=\mathcal{L}(f_{i}(\bm{x}),\bm{y}).$
These definitions are used to simplify our notations and better demonstrate our proofs. Then, we bring our key observations that are shown in the following theorem.
Theorem 1.
[…]
$S\sim\mathcal{N}(0,(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{1/2}\mathrm{Cov}(\mathbb{L}(f_{i}))(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{1/2}).$
Proof of Theorem 1 can be found in Appendix A.
Remark 1.
Theorem 1 suggests that the adversarial examples converge in $\mathcal{O}(\frac{1}{\sqrt{T}})$, while the difference between population loss and empirical loss converges in $\mathcal{O}(\frac{1}{T})$.
When we replace $\mathcal{L}$ with the cross-entropy loss in Theorem 1, we obtain Theorem 2 by some simple derivations.
Theorem 2.
Following the notation and assumptions in Theorem 1, when using the cross-entropy loss, we have:
[…]
Remark 2.
[…]
Theorem 2 gives us motivations for further exploring the effects of ensemble cardinality $T$ as empirical loss converges with a factor of $T$. This leads us to recovering the scaling laws on black-box adversarial attacks.
3.2 Scaling Laws on Image Classifiers
Figure 3: Scaling Laws over model ensemble in a black-box setting. We plot the x-axis on a base-2 logarithmic scale, and we measure both the attack success rate (ASR) and average cross entropy-loss of target models. The cardinality of the model ensembles is varied from $2^{0}$ to $2^{6}$, with models randomly selected for each ensemble. We fit the results of 4 target models by lines for better demonstration.
Before demonstrating scaling laws, we first introduce our experiment settings and hyperparameters.
Surrogate Models. To ensure the transferability and diversity of models, we select 64 models from Torchvision [66] and Timm [84] with different hyperparameters, architectures, and training datasets – Alexnet [51], Densenet [40], Efficientnet [76], Googlenet [74], Inception-V3 [75], Maxvit [80], Mnasnet [77], Mobilenet [39], Regnet [93], Resnet [37], Shufflenet [96], Squeezenet [43], Vgg [72], ViT [27], WideResnet [95], Beit [4], Ghostnet [36], Deit [78], LCnet [14, 85], Repvgg [20], Dpn [11], Mit [90], Bit [49], CvT [88], Twins-SVT [13], and NFNet [6]. The full list of our surrogate models can be found in Appendix B.
Target Models. We carefully choose models that are both widely known and are different from training ensembles in architecture and hyperparameters. We select 7 models as our target models, including Swin-Transformers [61], ConvNeXt [62], and ResNeXt [91].
Dataset. Following previous work [47], we randomly take 100 images from the NIPS17 dataset***https://www.kaggle.com/competitions/nips-2017-non-targeted-adversarial-attack. All images are resized into 224 $\times$ 224. We select 5 categories as target labels, including cat, car, ship, deer and bird.
Hyperparameters. We set the perturbation budget $\epsilon=8/255$ under $\ell_{\infty}$ norm. We take 40 iteration steps of SSA-CWA, while other hyperparameter settings are the same as in [9]. We set the loss function $\mathcal{L}$ to be the cross-entropy loss between the output logit $f(\bm{x})$ and the label $\bm{y}$.
Metrics. We leverage attack success rate (ASR) as our metric, which is (adversarial examples that induce successful attacks /// all adversarial examples).
Experiment Results. The results of ensemble attacks are shown in Fig. 3. From our observations, we empirically derive the scaling laws between the attack success rate (ASR) and ensemble cardinality $T$ as follows:
$\text{ASR}=\alpha\log T+C.$ (9)
It can be explained by the following points:
  • $C$ is influenced by various factors, such as hyperparameters, model architectures, and adversarial targets, and therefore takes different values under different conditions.
  • For a fixed adversarial target, $\alpha$ depends solely on the model itself. Models with the same architecture share the same $\alpha$, regardless of hyperparameter settings.
  • When $T$ is small, the measured ASR exhibits large variance due to different model selections, suggesting that our scaling laws tend to be a statistical result.
  • ASR acts as a good representation of the empirical loss $\mathbb{L}$ defined in Eq. (2). It can be observed from Fig. 3 that there exists strong symmetry between $\mathbb{L}$ and ASR.
  • The scaling laws become less useful as the ASR approaches 1, since ASR no longer reflects $\mathbb{L}$ at that point. This indicates that our observations hold only when the success rate is below the converging threshold.
Our experiment results clearly demonstrate the universality of scaling laws, independent of specific test models or adversarial targets. It is important to note that both our training and evaluation models are open-sourced and pretrained on ImageNet [18], which means they do not strictly satisfy the i.i.d. assumption discussed in Sec. 3.1. Therefore, the theorems derived in Sec. 3.1 should be viewed as a sign of convergence and are not directly related with the empirical result measured by our experiments. Another key improvement observed from our scaling laws is that the evaluation loss $\mathbb{L}$ decreases exponentially with the ensemble cardinality $T$, which is much tighter compared to bounds derived in previous works [41, 94].
3.3 Scaling Laws on Defended Models
Figure 4: Scaling Laws for defended models. While our adversarial examples exhibit high attack success rates across defenses, clear linear scaling laws can be observed from the fitted lines.
Figure 5: Scaling Laws for MLLMs. We measure the ASR of target MLLMs and fit the results with lines. Note that scaling laws still exist, except for the fact that results vary largely between models and targets.
To protect neural networks from attacks, various defense mechanisms have been proposed. Our scaling laws work well with these defense mechanisms, showing that these laws tend to be a general result across defenses.
Target Models. We select six defended models as target models, including 2 adversarially trained models – FGSMAT [52] and Ensemble AT (EnsAT) [79], and 4 input purification defenses – high-level representation guided denoiser (HGD) [56], bit depth reduction (BDR) [24], JPEG compression [16], and DiffPure [67]. We use the same surrogate models and algorithms as Sec. 3.2.
Experiment Results. As demonstrated in Fig. 4, we observe a clear scaling law in all these defended models, further demonstrating its universality. By ensembling a large number of normally trained surrogate models in Sec. 3.2, we improve the attack success rate of DiffPure [67] by more than 30%, adversarially trained models by more than 20%, and high-level guided denoiser [56] by more than 70%. This demonstrates that simply increasing the number of surrogate models can break state-of-the-art defenses like DiffPure [67] and adversarial training [65].
3.4 Scaling Laws on MLLMs
Multimodal Large Language Models (MLLMs) [58, 68, 29] have demonstrated impressive abilities in vision understanding, captioning, and reasoning. In this section, we show that our scaling laws persist in these MLLMs.
Target Models. We still leverage the adversarial examples generated against traditional image classifiers and select 3 popular VLMs – LLaVA-v1.5-7b [58], DeepSeek-VL-7b [17], and Phi-3-vision [29] as target models.
Evaluation Metrics. As MLLMs output text rather than logits, we adopt the definition of ASR from MultiTrust [98], which leverages GPT-4 for determining whether the targeted object is stated in the description. In addition to assessing the correctness of the image classification, it introduces prompts to evaluate the degree of distortion in the crafted image. If GPT determines that the image is highly corrupted, the attack is considered unsuccessful, even if the target model misclassifies the image. This metric ensures that our judgments of successful attacks are of high quality.
Experiment Results. From the results presented in Fig. 5, several key observations can be drawn:
  • All models exhibit a similar value of $C$, indicating that they are all highly robust to attack patterns when the ensemble cardinality $T$ is small.
  • $\alpha$ varies significantly across models, revealing the considerable differences in robustness between MLLMs.
  • Since MLLMs are trained on multimodal data, different adversarial targets significantly influence the attack success rates. This variability may stem from the inherent characteristics of the target labels, as naturally some labels (such as “deer”) are less representative compared to labels like “cat”. This phenomenon, often referred to as class imbalance, has been discussed in prior studies [46].
When compared to image classifiers, MLLMs demonstrate superior robustness across all chosen adversarial targets. This supports the general opinion that multimodal training leads to more robust representations [42], enabling the model to focus more on the essential features of the images rather than specific attack patterns. Furthermore, in contrast to traditional views that scaling model parameters results in increased robustness, we find that among the three target models, Phi-3-vision exhibits the highest robustness. This suggests that the scaling laws derived previously in [5] may not be directly applicable to multimodal models.
4 Extending Scaling Laws to Vision Encoders
Figure 6: Scaling Laws when we use CLIPs as surrogate models under $\epsilon=8/255$. We take ASR measured by Metric 1 to ensure our results are of high confidence.
In this section, we extend our findings of scaling laws from image classifiers to multimodal foundation models such as CLIP [70]. Sec. 4.1 shows the extension of Eq. (2) to image embeddings, while Sec. 4.2 demonstrates high success rates over commercial multimodal large language models (MLLMs) under our framework.
4.1 Attacking Image Embeddings
Contrastive Language-Image Pretraining (CLIP) is a popular self-supervised framework that can pretrain large-scale language-vision models on web-scale text-image pairs via contrastive learning [70]. Compared to image classifiers that are normally trained on image classification datasets, CLIPs have demonstrated superior zero shot generalization abilities in multimodal downstream tasks [69, 53].
Based on the generalization ability of CLIP models, previous works have utilized them as surrogate models in model ensembles [26, 1, 32]. In fact, the targeted attack framework described in Eq. (2) can be easily extended by replacing the loss function $\mathcal{L}$ with the similarity measure function $\operatorname{sim}(\cdot)$ from CLIP (usually cosine embedding loss in practice), and by substituting the image classifiers $f$ with CLIP’s image encoders $f^{I}$. This modification leads to the following optimization objective:
$\min_{\bm{x}}\frac{1}{T}\sum_{j=1}^{T}\operatorname{sim}(f^{I}_{j}(\bm{x}),f^{I}_{j}(\bm{x}_{\text{tar}})),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (10)
Note that besides an empirical loss function, Eq. (10) can also be explained as minimizing the distance between the embeddings of the adversarial examples and the target images on a low-dimension embedding manifold.
4.2 Experiments
We first introduce our settings for attacking pretrained CLIP models and MLLMs used for our evaluation.
Surrogate Models. We select 12 pretrained open-source CLIP models via OpenCLIP [12]. In contrast to previous works that typically use fewer than 4 surrogate models, our ensemble choice is sufficient to yield robust results. The full list of our surrogate clips can be found in Appendix. E.
Target Models. We select 6 commercial MLLMs as our black-box test models - GPT-4o, GPT-4 Turbo [68], Claude-3.5 Sonnet https://www.anthropic.com/news/claude-3-family, Gemini-1.5 Pro https://deepmind.google/technologies/gemini/pro/, Hunyuan [30], Qwen [28] - and 3 open-source MLLMs - LLama-3.2 90B §§§https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Pixtral Large https://mistral.ai/news/pixtral-large/, and DeepSeek-VL 7B [17].
ASR Target: Cat Target: Car Target: Ship Target: Deer Target: Bird
$\epsilon$ ($\ell_{\infty}$ norm) $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$
GPT-4o 0.58 0.99 0.22 0.52 0.57 0.91 0.40 0.89 0.56 0.97
GPT-4-Turbo 0.68 0.99 0.23 0.55 0.59 0.92 0.38 0.86 0.61 0.97
Claude-3.5 0.56 0.91 0.18 0.42 0.41 0.79 0.37 0.74 0.38 0.80
Gemini-1.5 Pro 0.4 0.92 0.16 0.50 0.35 0.82 0.29 0.82 0.38 0.83
Hunyuan-Vision 0.58 0.95 0.23 0.53 0.59 0.90 0.42 0.88 0.52 0.95
Qwen-VL Plus 0.75 0.98 0.28 0.54 0.65 0.91 0.52 0.92 0.69 0.97
Llama-3.2 90B 0.48 0.88 0.21 0.48 0.42 0.87 0.39 0.87 0.4 0.86
Pixtral Large 0.49 0.88 0.19 0.49 0.46 0.87 0.36 0.77 0.44 0.87
Deepseek-VL 7B 0.54 0.96 0.17 0.52 0.52 0.87 0.44 0.91 0.51 0.93
Table 2: ASR under Metric 2 measured across different settings of $\epsilon$. Remarkably, combined with our methods, most adversarial examples are considered to contain the targets’ features by MLLMs given $\epsilon=16/255$.
Hyperparameters. To demonstrate the scaling laws, we set $\epsilon=8/255$. For practical applications, we follow the common settings in [26, 64] and set $\epsilon=16/255$. All other experiment settings are consistent with those in Sec. 3.
Evaluation Metrics. We employ two types of evaluation metrics in this study. The first metric is the same as in Sec. 3.4. The second metric focuses on the features of adversarial examples, whose goal is to evaluate the insertion of features within adversarial targets into these examples. It requires the target model to generate a description and asks GPT whether the description mentions the adversarial target. Both metrics take ASR as the final score, which we refer to as Metric 1 and Metric 2 in the following sections for simplicity. Test prompts can be found in Appendix D.
Experiment Results. Fig. 6 presents the experiment results on the test models under Metric 1. These results reveal that scaling laws continue to hold when using CLIP models as surrogate models. This outcome seems natural as image features are still correctly learned within the vision encoders. Our method achieves success rates of over 30% on most adversarial targets under Metric 1, indicating that more than $1/3$ of the original images are transformed into high-quality adversarial examples with minimal noise, which demonstrates the efficiency of our method.
We also conduct extensive experiments under Metric 2. From Table. 2 we observe that the high attack success rates of our adversarial examples are generalized results, and the transferability of our adversarial examples is strong enough to fool most popular MLLMs. Compared to previous methods that tried to attack commercial MLLMs by visual adversarial examples [26], our methods exhibit a 40% higher success rate. This demonstrates the strength of our core inspiration of scaling, and highlights the practical applications of our methods. Meanwhile, we find the recently released open-source models Llama-3.2 90B and Pixtral Large exhibiting a better performance than other models generally, which indicates that the AI community is running towards developing more robust, safe and reliable MLLMs.
5 Discussions
In this section, we further demonstrate the scaling laws by using different optimizers, visualizing the adversarial perturbations, and collecting the fitting error.
5.1 Scaling Laws are General Principles
The idea that enlarging the model ensemble induces higher attack success rates has been proposed in previous work [94]. Nevertheless, while prior work usually provides qualitative guesses or bounds, we exploit linear scaling laws by extensive quantitative experiments. While our scaling laws work well with targeted attack settings and the SSA-CWA algorithm, further experiment results reveal that our scaling laws are general principles.
(a) Under untargeted setting.
(b) Across different algorithms.
Figure 7: Ablations on untargeted settings and other algorithms. We mainly follow the experiment configurations in Sec. 3.1. For untargeted attacks, we set $\epsilon=2/255$. For other algorithms, we leverage cat as the target class and plot the average ASR over all target models. (a) It demonstrates our linear scaling laws work under untargeted settings. (b) It supports our claim that linear trends occur on variance-reduction algorithms, while MI-FGSM exhibits inferior performance due to gradient conflicts.
Untargeted Attack Setting. Untargeted attacks are similar to targeted attacks except for a different training objective from Eq. (2), and all our algorithms are obviously compatible with the untargeted setting:
$\max_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}_{\text{nat}}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (11)
Our experiment results in Fig. 7(a) serve as a clear sign that scaling laws remain valid over untargeted settings. We mainly follow our settings in Sec. 3.1, while we set $\epsilon=2/255$, and measure the results on 3 target models. Taking an extremely small $\epsilon$ is necessary for demonstrating our linear scaling laws, otherwise the attack success rate will easily reach $100\%$ and therefore come to a plateau. The results validate our claim that scaling laws exist regardless of specific attack targets.
Across Different Attack Algorithms. Besides SSA-CWA, we also provide experimental validations of different attack algorithms in Fig. 7(b). As stated in Sec. 2.2, severe gradient conflicts can be observed when we iteratively update the input image $\bm{x}$ if we directly perform backward propagation over the sum of output logits (e.g., as in MI-FGSM). Gradient conflicts prevent the input $\bm{x}$ from being updated properly and lead to missing information of model weaknesses. However, advanced algorithms, such as SVRE [92] and CWA [9] address this issue by leveraging optimizers that can reduce the variance of gradients. Our experiment results support our claim that scaling laws occur only on such algorithms. Note that the attack success rate of MI-FGSM also becomes higher with increasing ensemble cardinality, further validating that our theoretical insight from Theorem 2 is regardless of specific attack methods.
5.2 Visualizations of Adversarial Perturbations
Figure 8: Visualization of adversarial perturbations. A clear trend of the increasing semantic interpretability of adversarial perturbations over different target objects.
As shown in Fig. 8, the semantic meaning of our adversarial perturbations also increases as the cardinality of the model ensemble grows. This suggests that the attack success rates are closely related to the interpretability of the perturbations, aligning with human observations and preferences [15]. Since the ultimate goal of adversarial attacks is to generate high-quality adversarial examples, it is preferable to have interpretable perturbations, as this reinforces the idea that the scaling effectively helps learn common features rather than merely converging to local optima.
As people traditionally believe images lie on a low dimensional manifold of desirable properties [31], this interpretability of adversarial examples can be also understood within the feature space of images, where human-designed attack algorithms aim to eliminate the non-robust features of images and shift the image towards the robust features of the target class [45]. From this perspective, our approach of enlarging model ensembles seeks to provide more information about the manifold itself, offering more accurate representations of the image features.
When we enlarge $\epsilon$ from $8/255$ to $16/255$ or even larger, the effect of the natural image goes down to 0, and the adversarial attack task degenerates to an image generation task. We claim that our method of scaling model ensemble still performs well under the degenerated image generation task, where the visualized results are shown in Appendix F.
5.3 Fitting Error of Scaling Law Curve
To comprehensively demonstrate the universality of our scaling laws, we provide the MSE fitting error from Fig. 3 in Table.3. It can be observed that the small fitting error acts as a quantitative validation of our scaling laws. Thus, our scaling laws remain a scalable way to enhance the transferability of adversarial examples.
Targets Cat Car Ship Deer Bird
Metric ASR Loss ASR Loss ASR Loss ASR Loss ASR Loss
Fitting Error 0.0040 0.34 0.0040 0.17 0.0038 0.46 0.0025 0.05 0.0023 0.27
Table 3: The averaged MSE fitting error of our curves in Fig. 3. Note that ASR usually has range $[0,1]$, while loss has range $[0,10]$. The fitting error is averaged on all test models, not only the test models that are fitted with lines for demonstration.
6 Conclusion
In this work, we investigate the scaling laws for black-box adversarial attacks. Theoretically, we provide asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec. 3.1. Empirically, we show that the attack success rate and the loss are converging exponentially. We verify this by comprehensive experiments over traditional image classifiers (Sec. 3.2), defended models like AT and DiffPure (Sec. 3.3), and modern MLLMs (Sec. 3.4 and Sec. 4) using various optimizers (Sec. 5). These experimental results demonstrate the universality of scaling laws in transfer attacks.
Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than focusing on designing intricate algorithms on a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy multimodal large foundation models.
Appendix A Proof of Theorem 1.
As defined in Definition 1, $\bm{x}^{*}$ is the minimizer of the population loss $\mathbb{L}(\bm{x})=\mathbb{E}_{f\in\mathcal{F}}[\mathcal{L}(f(\bm{x}),\bm{y})]$, and $\hat{\bm{x}}$ is the minimizer of the empirical loss $\hat{\mathbb{L}}(\bm{x})=\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y})$. Thereby we have by gradient equals 0:
$\nabla\mathbb{L}(x^{*})=\nabla\hat{\mathbb{L}}(\hat{\bm{x}})=0$
We use Taylor expansion of $\nabla\hat{\mathbb{L}}$ at $\hat{\bm{x}}$:
$0=\nabla\hat{\mathbb{L}}(\hat{\bm{x}})=\nabla\hat{\mathbb{L}}(\bm{x}^{*})+\nabla^{2}\hat{\mathbb{L}}(\bm{x}^{*})(\hat{\bm{x}}-\bm{x}^{*})+\mathcal{O}(\|\hat{\bm{x}}-\bm{x}^{*}\|^{2}).$
Rearranging, we have:
$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\approx-\sqrt{T}(\nabla^{2}\hat{\mathbb{L}}(\bm{x}^{*}))^{-1}\nabla\hat{\mathbb{L}}(\bm{x}^{*}).$
From Central Limit Theorem:
$\sqrt{T}\nabla\hat{\mathbb{L}}(\bm{x}^{*})$ $=\sqrt{T}(\nabla\hat{\mathbb{L}}(\bm{x}^{*})-\nabla\mathbb{L}(\bm{x}^{*}))$
[…]
and the law of large numbers:
[…]
Plug into the previous equation using $A\mathcal{N}(0,\Sigma)=\mathcal{N}(0,A\Sigma A^{T})$:
[…]
The above indicate $\hat{\bm{x}}-\bm{x}^{*}$ converges in $\mathcal{O}(\frac{1}{\sqrt{T}})$.
We can also use Taylor expansion of $\mathbb{L}$ at $\bm{x}^{*}$:
$\mathbb{L}(\hat{\bm{x}})\approx\mathbb{L}(\bm{x}^{*})+0+\frac{1}{2}(\hat{\bm{x}}-\bm{x}^{*})^{T}\nabla^{2}\mathbb{L}(\bm{x}^{*})(\hat{\bm{x}}-\bm{x}^{*}).$
Rearranging, we have:
$T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))$ $\approx\frac{1}{2}\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})^{T}\nabla^{2}\mathbb{L}(\bm{x}^{*})\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})$
$=\frac{1}{2}\|\nabla^{2}\mathbb{L}(\bm{x}^{*})\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\|_{2}^{2}=\frac{1}{2}\|S\|_{2}^{2},$
where
$S\sim\mathcal{N}(0,\nabla^{2}\mathbb{L}(\bm{x}^{*})^{-1/2}\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*}))\nabla^{2}\mathbb{L}(\bm{x}^{*})^{-1/2}).$
Thus $T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))$ converges into a chi-square distribution, which means $\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*})$ converges in $\mathcal{O}(\frac{1}{T})$.
Appendix B Experiment Settings for Scaling Laws
In this section we list our models that are included in Sec. 3.2. We select these models based on their diversity in architectures, training methods, and training datasets. Results are shown in Table 4 and Table 5, where all of our models come from open-source platforms such as Torchvision [66] and Timm [84]. Note that default in the table means normally trained on ImageNet-1K dataset.
Model name Architecture Settings
AlexNet alexnet default
DenseNet densenet121 default
densenet161 default
densenet169 default
EfficientNet efficientnet-b0 default
efficientnet-b1 default
efficientnet-b4 default
efficientnet-v2-s default
efficientnet-v2-m default
efficientnet-v2-l default
GoogleNet googlenet default
Inception inception-v3 default
MaxVit maxvit-t default
MnasNet mnasnet-0-5 default
mnasnet-0-75 default
mnasnet-1-0 default
mnasnet-1-3 default
MobileNet mobilenet-v2 default
mobilenet-v3-s default
mobilenet-v3-l default
RegNet regnet-x-8gf default
regnet-x-32gf default
regnet-x-400mf default
regnet-x-800mf default
regnet-y-16gf default
regnet-y-32gf default
regnet-y-400mf default
regnet-y-800mf default
ResNet resnet18 default
resnet34 default
resnet50 default
resnet101 default
resnet152 default
ShuffleNet shufflenet-v2-x0.5 default
shufflenet-v2-x1.0 default
shufflenet-v2-x1.5 default
shufflenet-v2-x2.0 default
SqueezeNet squeezenet-1.0 default
squeezenet-1.1 default
ViT vit-b-16 default
vit-b-32 default
vit-l-16 default
vit-l-32 default
vit-b-16 pretrained on ImageNet-21K
Table 4: List of pretrained image classifiers that are selected for our model ensemble.
Model name Architecture Settings
Vgg vgg11 default
vgg11-bn default
vgg13 default
vgg13-bn default
vgg16 default
vgg16-bn default
vgg19 default
Wide-ResNet wide-resnet50-2 default
wide-resnet101-2 default
BeiT beit-B-16 pretrained on ImageNet-21k
GhostNet ghostnet-100 default
DeiT deit-B-16 default
LCnet lcnet-050 default
RepVgg repvgg-a2 default
DPN dpn98 default
SegFormer mit-b0 default
BiT bit-50 default
CvT cvt-13 default
Twins-SVT twins-svt-large default
NFNet nfnet-l0 default
Table 5: List of pretrained image classifiers that are selected for our model ensemble (continued).
Appendix C SSA-CWA Algorithm Declaration
We provide the pseudo-code version of SSA-CWA implementation in Algorithm 1. Their combination is induced from  [63] and  [9], while implementations of other algorithms (e.g., MI-FGSM, MI-CWA) can be found in their original papers.
Algorithm 1 SSA-CWA Algorithm
1:Input: natural image $\bm{x}_{\text{nat}}$, label $\bm{y}$, $L_{\infty}$ constraint $\epsilon$, iterations $T$, loss function $\mathcal{L}$, spectrum transformation function $\mathcal{T}$, model ensemble $\{f_{i}\}_{i=1}^{N}$, spectrum transformation times $M$, decay factor $\mu$, step sizes $\alpha,\beta$.
2:Output: An adversarial example $\bm{x}$
3:for t = 0 to T-1 do
4:     Update $\bm{x_{t}}$ by $\bm{x}^{0}_{t}=x_{t-1}$
5:     for i = 1 todo
6:         for j = 1 todo $\triangleright$ Spectrum transformation
7:              Obtain transformed image $\mathcal{T}(\bm{x}^{j}_{t,i})$
8:              Compute $\bm{g}_{i}^{j}=\nabla_{\bm{x}^{j}_{t,i}}\mathcal{L}(\mathcal{T}(\bm{x}^{j}_{t,i}),\bm{y})$
9:         end for
10:         Compute gradient $\bm{g}_{i}=\frac{1}{M}\sum_{j=1}^{M}\bm{g}^{j}_{i}$
11:         Update inner momentum by $\bm{m}=\mu\cdot\bm{m}+\frac{\bm{g}_{i}}{\left\lVert\bm{g}_{i}\right\rVert_{2}}$
12:         Update $\bm{x}^{i}_{t}=\text{clip}(\bm{x}_{\text{nat}},\epsilon)\{\bm{x}^{i-1}_{t}+\beta\cdot\bm{m}\}$
13:     end for
14:     Compute gradient $\bm{g}=\bm{x}_{t}^{N}-\bm{x}_{t}^{0}$
15:     Update outer momentum by $\bm{h}=\mu\cdot\bm{h}+\frac{\bm{g}}{\left\lVert\bm{g}\right\rVert_{2}}$
16:     $\bm{x}_{t}=\text{clip}(\bm{x}_{\text{nat}},\epsilon)\{\bm{x}_{t}^{N}+\alpha\cdot\text{sign}(\bm{h})\}$
17:end for
18:Return: $\bm{x}_{T-1}$
Appendix D Test Prompts as Metrics
As VLMs and MLLMs output prompts rather than logits, we evaluate the attack efficacy by mainly focusing on the “classification” results of the test models, which is leveraging a GPT4 assistant to judge the content of the test models’ descriptions. We develop Metric 1 and Metric 2, focusing on different aspects of the descriptions.
Fig. 9(a) and Fig. 9(b) display our prompts for Metric 1 and Metric 2 respectively. Metric 1 follows the settings in previous works [98], while Metric 2 is our customized metric to capture image features, as stated in Sec. 4.2. By comparing the results in Fig. 6 and Table. 2, Metric 2 have a 10% higher success rate than Metric 1 under $\epsilon=8/255$ cases. This phenomenon is natural since some images of high noise will be still considered as success in Metric 2, which is more aligned with our practical approaches [26].
(a) Prompts for Metric 1.
(b) Prompt for Metric 2.
Figure 9: Prompts of two metrics. (a) Besides prompt to test the adversarial example contain a target label, we also use a prompt to judge the image quality, ensuring our successful attacks are all of small noise and highly transferable. (b) The prompt concentrates on semantic meanings of the image descriptions. To consider the attack successful, it only needs the target to appear in the description.
Appendix E Experiment Settings for Attacking MLLMs
Model Name Architecture Datasets
ViT-L OpenAI ViT-L-14 WebImageText
ViT-B LAION-2B ViT-B-32 LAION-2B
ViT-B LAION-2B ViT-B-16 LAION-2B
ViT-bigG LAION-2B ViT-bigG-14 LAION-2B
ViT-g LAION-2B ViT-g-14 LAION-2B
ViT-H LAION-2B ViT-H-14 LAION-2B
ViT-L LAION-2B ViT-L-14 LAION-2B
ConvNeXt-L LAION-2B ConvNeXt-large-d-320 LAION-2B
ViT-B DataComp ViT-B-16 DataComp-1B
ViT-B DataComp ViT-B-32 DataComp-1B
ViT-L DataComp ViT-L-14 DataComp-1B
SigLip WebLI ViT-SO400M-14-SigLIP WebLI
Table 6: List of CLIPs that are selected as model ensemble.
Figure 10: Visualizations of adversarial examples under $\epsilon=64/255$. The adversarial examples capture basic information of the target class, such as the fur of cats, tires of vehicles, cargo on ships, antlers of deer, and feathers of birds.
In Table. 6 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [70], OpenCLIP [12], and Timm [84].
Appendix F Degenerated Image Generation Task
As stated in Sec. 5.2, when $\epsilon$ goes larger, the adversarial attack task degenerates to an image generation task. Previous works [9, 1, 82] do not ever think about this field, but we believe that adversarial attack task and image generation task are indeed homogenous. We display our adversarial examples generated from 64 image classifiers (stated in Appendix B) under $\epsilon=64/255$ in Fig. 10. From the visualizations we observe that our methods can capture the key features of the target class and reproduce them based on the original textures of images. This visualization further supports our methods are really attacking common weakness instead of running into local optima. We leave further explorations and analyses as future work.
Appendix G Related Work
G.1 Black-Box Attacks
Deep learning models are widely recognized for their vulnerability to adversarial attacks [73, 35]. Extensive research has documented this susceptibility, particularly in image classification contexts [2, 7], under two principal settings: white-box and black-box attacks. In the white-box setting, the adversary has complete access to the target model [65]. Conversely, in the black-box setting, such information is obscured from the attacker. With limited access to the target model, black-box attacks typically either construct queries and rely on feedback from the model [25, 44] or leverage the transferability of predefined surrogate models [60]. Query-based attacks often suffer from high time complexity and substantial computational costs, making transfer-based methods more practical and cost-effective for generating adversarial examples, as they do not necessitate direct interaction with target models during the training phase.
Within the realm of transfer-based attacks, existing methods can generally be categorized into three types: input transformation, gradient-based optimization, and ensemble attacks. Input transformation methods employ data augmentation techniques before feeding inputs into models, such as resizing and padding [89], translations [23], or transforming them into frequency domain [63]. Gradient-based methods concentrate on designing optimization algorithms to improve transferability, incorporating strategies such as momentum optimization [22, 57], scheduled step sizes [33], and gradient variance reduction [81].
G.2 Ensemble Attacks
Inspired by model ensemble techniques in machine learning [87, 19, 34], researchers have developed sophisticated adversarial attack strategies utilizing sets of surrogate models. In forming an ensemble paradigm, practitioners often average over losses [22], predictions [60], or logits [22]. Additionally, advanced ensemble algorithms have been proposed to enhance adversarial transferability [55, 9, 92, 8]. However, these approaches tend to treat model ensemble as a mere technique for improving transferability without delving into the underlying principles.
Prior studies [41] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization. Recent research [94] has theoretically defined transferability error and sought to minimize it through model ensemble. Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [59].
G.3 Scaling Laws
In recent years, the concept of scaling laws has gained significant attention. These laws describe the relationship between model size – measured in terms of parameters, data, or training FLOPs – and generalization ability, typically reflected in evaluation loss [48]. Subsequent studies have expanded on scaling laws from various perspectives, including image classification [3, 38], language modeling [71] and mixture of experts (MoE) models [30], highlighting the fundamental nature of this empirical relationship.
In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [50, 5]. However, the broader concept of scaling remains underexplored. While some preliminary observations on the effect of enlarging model ensembles can be found in prior studies [94], these effects have not been systematically analyzed. Recent work Wei et al. [83] also explore the scaling law of attack success rate when jailbreaking large language models [99, 10] by increasing number of in-context demonstrations In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models. Motivated by this perspective, we unveil vulnerabilities within commercial LLMs by targeting their vision encoders.

### v4

@@
 Abstract
-Adversarial examples usually exhibit good cross-model transferability, enabling attacks on black-box models with limited information about their architectures and parameters, which are highly threatening in commercial black-box scenarios.
-Model ensembling is an effective strategy to improve the transferability of adversarial examples by attacking multiple surrogate models.
-However, since prior studies usually adopt few models in the ensemble, there remains an open question of whether scaling the number of models can further improve black-box attacks.
-Inspired by the scaling law of large foundation models, we investigate the scaling laws of black-box adversarial attacks in this work.
-Through theoretical analysis and empirical evaluations, we conclude with clear scaling laws that using more surrogate models enhances adversarial transferability.
-Comprehensive experiments verify the claims on standard image classifiers, diverse defended models and multimodal large language models using various adversarial attack methods.
-Specifically, by scaling laws, we achieve 90%+ transfer attack success rate on even proprietary models like GPT-4o.
-Further visualization indicates that scaling laws exists on the interpretability and semantics of adversarial perturbations.
+Adversarial examples exhibit cross-model transferability, enabling threatening black-box attacks on commercial models.
+Model ensembling, which attacks multiple surrogate models, is a known strategy to improve this transferability.
+However, prior studies typically use small, fixed ensembles, which leaves open an intriguing question of whether scaling the number of surrogate models can further improve black-box attacks.
+In this work, we conduct the first large-scale empirical study of this question.
+We show that by resolving gradient conflict with advanced optimizers, we discover a robust and universal log-linear scaling law through both theoretical analysis and empirical evaluations: the Attack Success Rate (ASR) scales linearly with the logarithm of the ensemble size $T$.
+We rigorously verify this law across standard classifiers, SOTA defenses, and MLLMs, and find that scaling distills robust, semantic features of the target class.
+Consequently, we apply this fundamental insight to benchmark SOTA MLLMs.
+This reveals both the attack’s devastating power and a clear robustness hierarchy: we achieve 80%+ transfer attack success rate on proprietary models like GPT-4o, while also highlighting the exceptional resilience of Claude-3.5-Sonnet.
+Our findings urge a shift in focus for robustness evaluation: from designing intricate algorithms on small ensembles to understanding the principled and powerful threat of scaling.
 22footnotetext: Our code is available at https://github.com/liuchuan22/ScaleUpAttack.
 1 Introduction
-Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread applications [54].
-However, deep neural networks (DNNs) remain susceptible to adversarial examples [35, 73] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions.
-Adversarial attacks have garnered considerable interest due to their implications for understanding DNN functionality [21], assessing model robustness [98], and informing the development of more secure algorithms [65].
+Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread applications [lecun2015deep].
+However, deep neural networks (DNNs) remain susceptible to adversarial examples [goodfellow2014explaining, szegedy2013intriguing] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions.
+Adversarial attacks have garnered considerable interest due to their implications for understanding DNN functionality [dong2017towards], assessing model robustness [zhang2024benchmarking], and informing the development of more secure algorithms [madry2017towards].
 Figure 1: Key observation of scaling laws for the transfer-based black-box attack.
-By scaling the number of surrogate models in the ensemble, we observe that the attack success rate and the discriminative loss on the target model follow a clear scaling law.
+By scaling the number of surrogate models in the ensemble, we observe that the attack success rate and the discriminative loss on the target model follow a log-linear scaling law.
 Adversarial attacks are typically categorized into white-box and black-box attacks based on whether the attacker has knowledge of the architecture and parameters of the target model.
-Given the greater risks with commercial services, e.g., GPT-4 [68], black-box attack methodologies [60, 22] have attracted growing attention.
+Given the greater risks with commercial services, e.g., GPT-4 [openai2024gpt4technicalreport], black-box attack methodologies [liu2016delving, dong2018boosting] have attracted growing attention.
 One prominent approach within black-box attacks is the transfer-based attack, which approximately optimizes the adversarial examples against the black-box model by minimizing the discriminative loss on a selection of predefined surrogate models.
-Researchers have devised advanced optimization algorithms [65, 81, 57] and data augmentation techniques [63, 23] in transfer-based attacks.
-Model ensembling [60, 22] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the transferability of adversarial examples.
-By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [22, 9].
-Despite the progress in ensemble attack algorithms, the number of surrogate models used in previous research is often limited [9, 97, 94].
-Recent advancement in large language models shows that scaling training data can significantly improve the generalization ability [48].
-Since the transfer-based attack is also a generalization problem that expects adversarial examples from surrogate models to generalize to target models, while the training data are surrogate models rather than text or images [22, 9], this equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attacks by scaling the number of surrogate models, akin to the scaling of data and model parameters in large language models like GPT?
-In this work, we investigate the scaling laws for black-box adversarial attacks.
-Theoretically, we provide an asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec.
-3.1.
-Empirically, we show that the attack success rate and the loss are converging exponentially.
-We verify this by comprehensive experiments over traditional image classifiers (Sec.
-3.2), defended models including adversarial training [65, 86] and DiffPure [67] (Sec.
-3.3), and modern multimodal large language models (Sec.
-3.4 and Sec.
-4) using various attack methods (Sec.
-5).
-These experimental results demonstrate the universality of scaling laws in transfer-based attacks.
-More concretely, through scaling laws, we achieve over 30% improvement against vision classifiers and over 10% improvement against defended models.
-By extending the surrogate models to a group of CLIP [70] vision encoders, we achieve over 90% attack success rate against GPT-4o and GPT-4-Turbo, and over 80% against other commercial MLLMs.
-This highlights the vulnerabilities of these models to perturbations and the significance of scaling laws for ensemble attacks.
-We also demonstrate clear scaling laws on the interpretability and semantics of adversarial perturbations.
-Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than designing intricate algorithms for a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy MLLMs.
+Researchers have devised advanced optimization algorithms [madry2017towards, wang2021enhancing, lin2019nesterov] and data augmentation techniques [long2022frequency, dong2019evading] in transfer-based attacks.
+Model ensembling [liu2016delving, dong2018boosting] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the transferability of adversarial examples.
+By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [dong2018boosting, chen2023rethinking].
+Despite the progress in ensemble attack algorithms, the number of surrogate models used in previous research is often limited [chen2023rethinking, zhang2024does, yao2025understandingmodelensembletransferable].
+Recent advancement in large language models shows that scaling training data can significantly improve the generalization ability [kaplan2020scalinglawsneurallanguage].
+Since the transfer-based attack is also a generalization problem that expects adversarial examples from surrogate models to generalize to target models, while the training data are surrogate models rather than text or images [dong2018boosting, chen2023rethinking], this equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attacks by scaling the number of surrogate models, akin to the scaling of data and model parameters in large language models like GPT?
+In this work, we conduct the first large-scale empirical investigation into the scaling laws of black-box adversarial attacks, given that the gradient conflict issue is properly resolved (Sec.
+2.2).
+Our central finding is the discovery of a robust and universal log-linear scaling law: the Attack Success Rate (ASR) on black-box targets scales linearly with the logarithm of the ensemble cardinality $T$.
+We provide an asymptotic analysis on the convergence rate of generation error and adversarial examples under idealized IID assumptions (Sec.
+3.1).
+Our core contribution, however, is the empirical discovery and verification of the precise log-linear form of these scaling laws (Sec.
+3.3).
+We rigorously demonstrate the universality of these scaling laws across diverse target families (Sec.
+3.4), and provide a qualitative explanation for this phenomenon (Sec.
+5.2): as $T$ increases, the optimizer is forced to discard model-specific, non-robust features and instead learns the robust, semantic features of the target class, resulting in highly transferable perturbations.
+Inspired by this fundamental insight, we then extend our scaling laws to attack the foundational vision encoders (e..g., CLIP) of modern MLLMs (Sec.
+4) to demonstrate its practical power as both a SOTA attack and a novel robustness benchmark.
+We achieve devastatingly high ASR on top-tier models, including 90.0% ASR on Qwen3-VL-235B and 85.0% ASR on GPT-4o.
+Meanwhile, our scaling benchmark reveals a stark hierarchy in SOTA model robustness.
+We highlight the exceptional resilience of Claude-3.5-Sonnet in sharp contrast to the high vulnerability of Gemini-2.5-Pro.
+These findings show that scaling is a critical, measurable threat vector, and we urge the community to adopt large-scale robustness evaluation, rather than designing intricate algorithms on small ensembles, to build and validate genuinely robust and trustworthy foundation models.
 2 Backgrounds
-In this section, we provide necessary backgrounds, including problem formulation and algorithms.
-2.1 Problem Formulation
-In this work, we focus on the challenging targeted attack setting.
+In this section, we provide necessary backgrounds, including the problem formulations for different attack settings and the prerequisite algorithms.
+2.1 Problem Formulations
+Our work investigates scaling laws across multiple black-box attack scenarios. We formulate the settings below.
+2.1.1 Attacks on Image Classifiers
 Let $\mathcal{F}$ denote the set of all image classifiers.
-Given a natural image $\bm{x}_{\text{nat}}$ and target class $\bm{y}$, the goal of targeted transfer-based attacks is to craft an adversarial example $\bm{x}$ that is misclassified to $\bm{y}$ by all models in $\mathcal{F}$.
-It can be formulated as a constrained optimization problem:
-$\min_{\bm{x}}\mathbb{E}_{f\in\mathcal{F}}\left[\mathcal{L}(f(\bm{x}),\bm{y})\right],\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (1)
-where $\mathcal{L}$ is the loss function (e.g., cross-entropy loss), and we consider adversarial perturbations in $\ell_{\infty}$ norm.
-Generally, we take a limited set of “training” classifiers $\mathcal{F}_{T}:=\{f_{i}\}_{i=1}^{T}\subset\mathcal{F}$ as model ensemble to approximate problem (1) and therefore the optimization objective becomes:
-$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (2)
-where $T=\left\lvert\mathcal{F}_{T}\right\rvert$ is the cardinality of model ensemble.
+The vulnerability of such classifiers to adversarial examples has been a long-standing area of research [szegedy2013intriguing, goodfellow2014explaining].
+Given a natural image $\bm{x}_{\text{nat}}$ with its true label $\bm{y}_{\text{nat}}$, the goal of a transfer-based attack is to craft an adversarial example $\bm{x}$ within an $\ell_{\infty}$ perturbation budget $\epsilon$ that fools black-box target models.
+Targeted Attack.
+The primary focus of our work is the challenging targeted attack setting.
+Given a target class $\bm{y}$ (where $\bm{y}\neq\bm{y}_{\text{nat}}$), the goal is to craft an $\bm{x}$ that is misclassified as $\bm{y}$.
+This can be formulated as minimizing the expected loss over all models in $\mathcal{F}$:
+$\min_{\bm{x}}\mathbb{E}_{f\in\mathcal{F}}\left[\mathcal{L}(f(\bm{x}),\bm{y})\right],\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (1)
+where $\mathcal{L}$ is a loss function (e.g., cross-entropy). In practice, we approximate this by ensembling a limited set of $T$ surrogate classifiers $\{f_{i}\}_{i=1}^{T}\subset\mathcal{F}$:
+$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}),\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (2)
+Untargeted Attack.
+We also validate our scaling laws in the untargeted setting in Sec.
+5.1, where the goal is to mislead the classifier to any class other than $\bm{y}_{\text{nat}}$.
+The objective is to maximize the loss with respect to original $\bm{y}_{\text{nat}}$:
+$\max_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}_{\text{nat}}),\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (3)
+2.1.2 Attacks on Vision Encoders
+Contrastive Language-Image Pretraining (CLIP) is a popular self-supervised framework that can pretrain large-scale language-vision models on web-scale text-image pairs via contrastive learning [radford2021learningtransferablevisualmodels, palatucci2009zero, lampert2009learning].
+Based on the generalization ability of CLIP models, previous works have utilized them as surrogate models in model ensembles [dong2023robust, anonymous2024tuap, fort2024ensembleeverywheremultiscaleaggregation].
+We extend our analysis to attack these vision encoders, which are foundational to modern MLLMs.
+Instead of classifying to a specific logit, the goal is to craft an image $\bm{x}$ whose embedding $f^{I}(\bm{x})$ is close to the text embedding $\bm{e}_{\text{tar}}$ of a target label $\bm{y}$.
+This target embedding is typically pre-computed using a text encoder, $f^{T}$.
+The objective function is adapted by replacing the classifier $f_{i}$ with a surrogate image encoder $f_{i}^{I}$ and the cross-entropy loss $\mathcal{L}$ with a similarity-based loss $\mathcal{L}_{\text{sim}}$ (e.g., cosine embedding loss):
+$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}_{\text{sim}}(f_{i}^{I}(\bm{x}),\bm{e}_{\text{tar}}),\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (4)
+where $\bm{e}_{\text{tar}}=f^{T}(\bm{y})$ is the fixed target text embedding.
 2.2 Prerequisite Algorithms
-Researchers have been developing sophisticated algorithms for ensemble attacks, and we select the state-of-the-art algorithms from them to conduct our experiments.
-MI-FGSM [22] is a pioneering method of transfer-based attacks. It stabilizes the gradient update directions by integrating a momentum term into the iteration:
+Researchers have developed sophisticated algorithms to enhance adversarial transferability, which can be categorized into gradient optimizers, input transformations, and ensemble strategies.
+2.2.1 Gradient Optimizers
+These methods refine the gradient update step itself. A pioneering example is MI-FGSM [dong2018boosting], which stabilizes gradients by integrating a momentum term:
 $\bm{g}_{t+1}=\mu\cdot\bm{g}_{t}+\frac{\nabla_{\bm{x}}\mathcal{L}(\bm{x_{t}},\bm{y})}{\left\lVert\nabla_{\bm{x}}\mathcal{L}(\bm{x_{t}},\bm{y})\right\rVert_{1}},$ (5)
-and updates $\bm{x}$ by $\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1})$.
-Common Weakness Method (CWA) [9] seeks to identify common vulnerabilities across surrogate models by encouraging flatter loss landscapes and closer proximity between the local optima of these models.
-It minimizes the second-order Taylor series:
+with update $\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1})$. Other methods like VMI-FGSM [wang2021enhancing] further refine this momentum concept.
+2.2.2 Input Transformations
+These methods apply data augmentation during the attack to find more generalizable adversarial patterns.
+Prominent examples include Diverse Inputs (DI) [xie2019improving] and Spectrum Simulation Attack (SSA) [long2022frequency].
+SSA, a state-of-the-art method, applies spectrum transformations $\mathcal{T}$ and computes gradients on the transformed input $\mathcal{T}(\bm{x})$.
+Figure 2: Comparisons between CWA and a naïve ensemble (MI-FGSM with logit averaging).
+As $T$ increases, the naïve method suffers optimization stagnation (gradient norm collapses) due to gradient conflict.
+Advanced methods like CWA, which align gradients, avoid this issue entirely.
+2.2.3 Ensemble Strategies and Gradient Conflict
+The above components can be applied to ensemble attacks.
+The most straightforward method, often used with MI-FGSM, is logit ensembling [dong2018boosting], where the optimizer computes gradients over the sum of all models’ logits.
+While simple, this strategy exposes a critical flaw: gradient conflict.
+We discover that as the number of surrogates ($T$) increases, the gradients from different models, which often lack semantics and have random directions, tend to cancel each other out.
+As shown in Fig.
+2, this conflict causes the ensemble gradient norm to collapse, leading to complete optimization stagnation.
+To address this, advanced optimizers are required.
+These methods aim to find a common direction of vulnerability rather than simply averaging gradients.
+They typically work by reducing gradient variance, such as SVRE [xiong2022stochastic], or by encouraging gradient alignment, which is the approach taken by the Common Weakness Attack (CWA) [chen2023rethinking].
+CWA seeks common vulnerabilities by encouraging flatter loss landscapes and minimizing the second-order Taylor series:
 $\frac{1}{T}\sum_{i=1}^{T}\left[\mathcal{L}(f_{i}(\bm{p}_{i},\bm{y})+\frac{1}{2}(\bm{x}-\bm{p}_{i})^{\top}H_{i}(\bm{x}-\bm{p}_{i})\right],$ (6)
-where $\bm{p}_{i}$ denotes the closest optimum of model $f_{i}$ to $\bm{x}$ and $H_{i}$ is the Hessian matrix at $\bm{p}_{i}$.
-Spectrum Simulation Attack (SSA) [63] applies a spectrum transformation $\mathcal{T}$ to the $\bm{x}$ and calculate gradients with respective to $\mathcal{T}(\bm{x})$ instead of $\bm{x}$, and can be combined with CWA to form SSA-CWA algorithm (pseudocode is provided in Appendix C).
-As shown in Table 1, SSA-CWA gives the best attack success rate.
-Therefore, in this paper, we use SSA-CWA as the core algorithm.
-We also explore the scaling laws of other algorithms in Sec.
-5.1.
-ASR Target: Car Target: Ship
-Methods MI-CWA SSA-CWA MI-CWA SSA-CWA
-Swin-v2 0.44 0.65 0.36 0.62
-ConvNeXt 0.52 0.58 0.53 0.62
-ResNeXt50 0.69 0.79 0.66 0.77
-Table 1: Comparison of attack success rate (ASR) between SSA and MI algorithms. We follow the experiment settings in Sec. 3.2, and measure the ASR on 2 target classes and 3 black-box models.
-Figure 2: Comparisons between CWA and MI-FGSM methods.
-We draw inputs $\bm{x}$ randomly from the dataset in Sec.
-3.2 and measure the averaged gradient magnitudes on $\bm{x}$ after 10 steps of iterations.
-The results are recorded across $T$ from $1$ to $30$.
-On the right side, there are the visualizations of gradients of $T=30$.
-Mitigating Gradient Conflict.
-We discover a new phenomenon of gradient conflict when ensembling a large number of surrogates.
-As shown in Fig.
-2, as the number of models increases, the conflict between gradients of different models using common optimizers (e.g., MI-FGSM) becomes more severe, eventually reducing the ensemble gradient to zero and resulting in complete optimization stagnation.
-We conjecture this is because the gradients of normally trained models commonly lack semantics and have random directions; thus, naïve ensemble methods that directly average these gradients lead to strong gradient conflicts, significantly impacting optimization.
-To address this problem, we suggest to use optimizers that reduce the variance of gradients (e.g., SVRE [92]) or encourage the cosine similarity of gradients (e.g., CWA [9]).
-As demonstrated in Fig.
-2, these methods fully avoid this issue.
-We analyze the gradient conflict on different attackers in more detail in Sec.
+where $\bm{p}_{i}$ is the closest optimum of model $f_{i}$ to $\bm{x}$ and $H_{i}$ is the Hessian. As Fig. 2 demonstrates, CWA effectively avoids the optimization stagnation seen in MI-FGSM.
+Our Chosen Algorithm: SSA-CWA.
+For our main experiments, we require an algorithm that both resolves the gradient conflict and achieves state-of-the-art performance.
+We therefore select SSA-CWA, which combines the advanced input transformation of SSA [long2022frequency] with the robust gradient optimization of CWA [chen2023rethinking] (pseudocode provided in Appendix C).
+While leveraging SSA-CWA as our core algorithm, we also explore the scaling laws of other optimizers and validate our gradient conflict hypothesis in Sec.
 5.1.
@@
 Sec.
-3.1 introduces fundamental concepts and theorems that give us motivations for the study of scaling laws.
+3.1 introduces the theoretical motivation for our study.
 Sec.
-3.2 contains our experiment settings and derivations of empirical scaling laws.
+3.2 details our experimental setup for generating adversarial examples.
 Sec.
-3.3 demonstrates clear scaling laws when attacking common defended models.
-Sec.
-3.4 shows the results over multimodal large language models.
-3.1 Motivations
+3.3 presents the initial discovery of the empirical scaling laws on standard classifiers.
+Finally, Sec.
+3.4 verifies the universality of these laws by testing them against state-of-the-art defended models and across diverse MLLMs.
+3.1 Motivation
 First, we observe that in most practical scenarios, $\bm{x}_{\text{nat}}$ and $\bm{y}$ are predefined and fixed.
@@
 Theorem 1.
-[…]
-[…]
-[…]
+Suppose that model ensemble $\{f_{i}\}_{i=1}^{T}$ is i.i.d.
+sampled from some distribution on $\mathcal{F}$, and $\hat{\bm{x}}$ goes to $\bm{x}^{*}$ as $T\to\infty$.
+Denote $\overset{d}{\to}$ as convergence in distribution, and $\mathrm{Cov}$ as the covariance matrix.
+Then we have the following asymptotic bound as $T\to\infty$:
+$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\overset{d}{\to}\mathcal{N}(0,(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}\mathrm{Cov}(\mathbb{L}(f_{i}))(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}),$ (9)
+$T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))\overset{d}{\to}\frac{1}{2}\|S\|_{2}^{2},\text{ with}$
 $S\sim\mathcal{N}(0,(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{1/2}\mathrm{Cov}(\mathbb{L}(f_{i}))(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{1/2}).$
@@
 Following the notation and assumptions in Theorem 1, when using the cross-entropy loss, we have:
-[…]
-[…]
+$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\overset{d}{\to}\mathcal{N}(0,(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}),$ (10)
+$T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))\overset{d}{\to}\frac{1}{2}\|\mathcal{N}(0,\bm{I})\|_{2}^{2}.$
 Remark 2.
-[…]
-Theorem 2 gives us motivations for further exploring the effects of ensemble cardinality $T$ as empirical loss converges with a factor of $T$.
-This leads us to recovering the scaling laws on black-box adversarial attacks.
-3.2 Scaling Laws on Image Classifiers
+Theorem 2 holds since $\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*})))=\nabla^{2}\mathbb{L}(\bm{x}^{*})$ for the cross-entropy loss.
+Therefore, we have $2T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))\overset{d}{\to}\chi^{2}(D)$ where $D$ is the dimension of the input, which means $\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*})$ converges in $\mathcal{O}(\frac{D}{2T})$.
+These theorems provide the core theoretical motivation for our work: increasing the ensemble cardinality $T$ is a principled approach, as it provably closes the gap between the empirical loss and the true population loss.
+However, we must acknowledge a crucial gap between theory and practical setting of black-box attacks.
+The theorems rely on i.i.d.
+assumptions that are not met in practice but for a theoretical convenience.
+In practice, our surrogate models are typically pre-trained on similar datasets (e.g., ImageNet) and share strong architectural biases, making them highly correlated [ilyas2019adversarialexamplesbugsfeatures].
+Consequently, the asymptotic $\mathcal{O}(1/T)$ convergence rate derived in Remark 2 is unlikely to hold.
+Therefore, we posit that while the theory provides a strong qualitative motivation (i.e., scaling $T$ should improve transferability), it does not offer a quantitative prediction for the empirical scaling law.
+The central question of how ASR and loss actually behave as $T$ increases in a practical setting remains open.
+This motivates our large-scale empirical investigation in the following sections to discover the true scaling laws governing black-box attacks.
+3.2 Experimental Setup
 Figure 3: Scaling Laws over model ensemble in a black-box setting.
@@
 The cardinality of the model ensembles is varied from $2^{0}$ to $2^{6}$, with models randomly selected for each ensemble.
-We fit the results of 4 target models by lines for better demonstration.
-Before demonstrating scaling laws, we first introduce our experiment settings and hyperparameters.
+We fit the results of selected target models by lines for better demonstration.
+To investigate the relationship between the number of surrogate models and attack efficacy, we first define our comprehensive experimental setup.
+This setup is used to generate a single, fixed set of adversarial examples, which are then evaluated across diverse target models to discover and verify our proposed scaling laws.
+3.2.1 Adversarial Example Generation
+Surrogate Pool.
+To ensure the transferability and diversity of our attack, we select a large pool of 64 models from Torchvision [torchvision2016] and Timm [rw2019timm].
+This pool includes a wide variety of architectures, hyperparameters, and training datasets, such as AlexNet [krizhevsky2012imagenet], DenseNet [huang2017densely], EfficientNet [tan2019efficientnet], ResNet [he2016deep], ViT [dosovitskiy2020image], DeiT [touvron2021training], and many others.
+The full list is available in Appendix B.
+Ensemble Sampling and Attack Algorithm.
+We generate adversarial examples by attacking ensembles of surrogate models.
+For each attack, an ensemble is constructed by randomly sampling $T$ models from our 64-model pool.
+We vary the ensemble cardinality $T$ across a logarithmic scale from $2^{0}$ to $2^{6}$.
+We use the SSA-CWA attack [chen2023rethinking] for 40 iteration steps, with a perturbation budget of $\epsilon=8/255$ under the $\ell_{\infty}$ norm.
+The loss function $\mathcal{L}$ is the standard cross-entropy loss.
+This process yields 7 distinct sets of adversarial examples, one for each value of $T$.
+3.2.2 Target Model Families
+To test the universality of any observed scaling phenomena, we evaluate our pre-generated adversarial examples against three distinct families of black-box target models:
+  • Family 1: Standard Image Classifiers. We select 7 widely-used models, including Swin-Transformers [liu2021swin], ConvNeXt [liu2022convnet], and ResNeXt [xie2017aggregated].
+  • Family 2: Defended Models.
+We select 6 models representing SOTA defenses, including adversarially trained models (FGSMAT [kurakin2016adversarial], EnsAT [tramer2017ensemble]) and purification defenses (HGD [liao2018defenseadversarialattacksusing], BDR [dong2020benchmarking], JPEG, and DiffPure [nie2022diffusionmodelsadversarialpurification]).
+  • Family 3: Multimodal Large Language Models.
+We select 3 popular MLLMs: LLaVA-1.5-7B [liu2023llava], Qwen-2.5-VL-7B [bai2025qwen25vltechnicalreport], and Llama-3.2-11B-Vision***https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/.
+These are used to further validate the generality of the laws across different data modalities and architectures.
+3.2.3 Dataset and Evaluation Metrics
+Dataset and Generation Protocol.
+Following previous work, we leverage the 1000 images from the ImageNet-1K NIPS 2017 dataset [nips-2017-non-targeted-adversarial-attack, yang2022boostingtransferabilitytargetedadversarial].
+To ensure the generality of our findings and create a large-scale testbed, we target 8 distinct categories for each source image [zhang2020understandingadversarialexamplesmutual].
+This creates a total of $1000\times 8=8000$ unique source-target pairs for each ensemble configuration.
+As we generate adversarial examples for 7 distinct ensemble cardinalities $T\in\{2^{0},2^{1},\dots,2^{6}\}$, this results in a large-scale dataset of $7\times 8000=56000$ generated adversarial examples.
+To ensure our results are statistically robust and not biased by a single ‘lucky’ or ‘unlucky’ random selection of surrogate models, we employ a batched generation protocol.
+For any given cardinality $T$, we first partition the 1000 source images into 8 disjoint batches of 125 images.
+Then, for each batch, we randomly sample a new ensemble of $T$ models from our surrogate pool.
+This newly sampled ensemble is used to craft the 125 adversarial examples (against all 8 target classes) for its assigned batch.
+This methodology ensures that the final metrics for each $T$ are averaged over 8 distinct, randomly-drawn ensembles, providing a highly robust and generalizable measurement of the scaling effect.
+Metrics.
+For Family 1 and 2, we measure the Attack Success Rate (ASR), which is defined by the ratio between number of successful attacks (misclassified as target label) and total number of attacks.
+We also record the cross entropy loss between classifiers’ predictions and target labels for Family 1 to demonstrate the validity of using ASR as the measurements of transferability.
+For Family 3 (MLLMs), which output text, we adopt a simple evaluation paradigm by string matching the target label text (e.g., ‘sea lion’ for target label 150) and the MLLMs’ descriptions.
+Since MLLMs produce descriptive text rather than precise classification results, we therefore adopt a relaxed keyword map (provided in Appendix D) to more robustly identify any semantic mentions of the target concept within the outputs.
+3.3 Log-Linear Scaling Laws in Image Classifiers
+3.3.1 Empirical Observation
+We begin our analysis by evaluating the generated adversarial examples on the standard image classifiers.
+As demonstrated in Fig.
+3 (top row), we identify a robust empirical phenomenon: the attack success rate (ASR) exhibits strong log-linear scaling laws with the ensemble cardinality $T$.
+As shown by the fitted lines, the ASR increases linearly as we increase $T$ on a logarithmic scale.
+Concurrently, the average cross-entropy loss against the target label (top row of Fig.
+3) exhibits a mirrored log-linear decay.
+This strong symmetric relationship confirms that ASR is a valid and robust proxy for the true optimization objective, as the decreasing loss directly correlates with the increasing success rate.
+This suggests that the effectiveness of ensemble attacks scales in a highly predictable manner, where each doubling of $T$ yields a consistent improvement in ASR.
+3.3.2 Analysis of the Log-Linear Relationship
+We empirically analyze the observable characteristics of these log-linear scaling laws:
+  • The Intercept (Baseline Vulnerability): The starting point of the line (at $T=1$) represents the model’s baseline vulnerability to single-model transfer attacks.
+As seen in Fig.
+3, this intercept is influenced by various factors, including the target model’s architecture and the specific adversarial target.
+  • The Slope (Scaling Efficiency): The slope of the fitted line represents the ‘scaling efficiency’ of the attack.
+It quantifies how much ASR is gained for each doubling of $T$.
+Our results suggest this slope is a key characteristic of the target model’s robustness to scaled attacks.
+  • Domain of Validity: Our observed scaling laws hold true within a specific domain.
+When $T$ is small ($2^{0},2^{1}$), the measured ASR shows large variance depending on the specific models randomly selected.
+The stable log-linear law emerges as a statistical result as $T$ increases.
+Meanwhile, as ASR approaches 1.0, the metric becomes saturated, and the linear trend naturally flattens.
+Our observations hold when ASR is below this saturation threshold.
+3.4 Verification of the Scaling Laws’ Universality
+(a) Scaling laws on Family 2.
+(b) Scaling laws on Family 3.
+Figure 4: Verification of the log-linear scaling laws’ universality.
+(a) The trend persists against SOTA defenses.
+(b) The trend persists for MLLMs.
+In both settings, we record the ASR averaged across all 8 target classes to show overall model robustness.
+Having established the existence of these log-linear scaling laws, we now test their universality.
+A critical question is whether this trend is a fragile artifact of standard classifiers or a more fundamental principle of transfer attacks.
+We test this by evaluating their robustness against active defenses (Family 2) and their generality across entirely different architectures and modalities (Family 3).
+3.4.1 Robustness Against Defended Models
+We first test the adversarial examples against the defended models (Family 2). The key question is whether SOTA defenses can ‘break’ this observed scaling trend.
+As demonstrated in Fig.
+4(a), the log-linear scaling laws remain robustly intact.
+While some defenses successfully lower the overall ASR (seen in the lower ASR values across all $T$), they do not eliminate the fundamental, predictable relationship between $\log T$ and ASR.
+This is a critical insight: defenses do not break the underlying scaling behavior.
+Instead, they primarily increase the ‘cost’ of a successful attack (i.e., requiring a larger $T$ to reach the same ASR).
+This finding demonstrates that scaling the ensemble is a viable path to breaking even SOTA defenses.
+By leveraging these scaling laws, we improved the ASR of DiffPure [nie2022diffusionmodelsadversarialpurification] by over 30%, adversarially trained models by over 20%, and HGD [liao2018defenseadversarialattacksusing] by over 50%.
+3.4.2 Generalization Across Architectures
+Finally, we perform the ultimate test of generality: do the adversarial examples, generated only from standard image classifiers, follow the same scaling laws on MLLMs?
+Remarkably, as shown in Fig.
+4(b), the log-linear scaling laws persist even in these advanced multimodal models.
+This cross-modal generalization is non-trivial, as the MLLMs possess vastly different architectures and are trained on different data modalities.
+This result strongly supports the generality of our scaling laws.
+While the log-linear form is general, the specific robustness of each model varies.
+For instance, LLaVA-1.5-7B and Qwen-2.5-VL-7B have similar parameter counts ($\approx 7$B) and both show high robustness to single-model attacks, yet their averaged ASR curves clearly diverge, indicating different vulnerabilities to scaling.
+This suggests that the relationship between model scale and adversarial robustness is more complex in MLLMs and challenges the simple ‘model scale = robustness’ narrative [bartoldson2024adversarialrobustnesslimitsscalinglaw].
+4 Extending Scaling Laws to Vision Encoders
+Figure 5: Verification of the log-linear scaling laws on SOTA MLLMs.
+We use an ensemble of CLIP models as surrogates, scaling the cardinality $T$ from 1 to 12.
+All attacks in this figure are generated using practical budget ($\epsilon=16/255$).
+The ASR is measured using our single, strict LLM-as-Judge metric (defined in Sec.
+4.1).
+The log-linear scaling laws clearly persist.
+In this section, we test the generality and power of our discovered scaling laws.
+Having established the log-linear scaling laws on standard classifiers, we now extend our investigation to a more complex and practical domain: attacking modern MLLMs by targeting their foundational vision encoders, such as CLIP [radford2021learningtransferablevisualmodels].
+We hypothesize that the same scaling principles hold when using the embedding-space attack formulation defined in Sec.
+2.1.2.
+4.1 Experimental Setup
+We first introduce our settings for attacking pretrained CLIP models and evaluating them on MLLM targets.
 Surrogate Models.
-To ensure the transferability and diversity of models, we select 64 models from Torchvision [66] and Timm [84] with different hyperparameters, architectures, and training datasets – Alexnet [51], Densenet [40], Efficientnet [76], Googlenet [74], Inception-V3 [75], Maxvit [80], Mnasnet [77], Mobilenet [39], Regnet [93], Resnet [37], Shufflenet [96], Squeezenet [43], Vgg [72], ViT [27], WideResnet [95], Beit [4], Ghostnet [36], Deit [78], LCnet [14, 85], Repvgg [20], Dpn [11], Mit [90], Bit [49], CvT [88], Twins-SVT [13], and NFNet [6].
-The full list of our surrogate models can be found in Appendix B.
+We select 12 pretrained open-source CLIP models via OpenCLIP [cherti2023reproducible].
+In contrast to previous works that typically use fewer than 4 surrogate models [zhao2023evaluatingadversarialrobustnesslarge], our scaled ensemble (up to $T=12$) is sufficient to yield robust results.
+The full model list is in Appendix E.
 Target Models.
-We carefully choose models that are both widely known and are different from training ensembles in architecture and hyperparameters.
-We select 7 models as our target models, including Swin-Transformers [61], ConvNeXt [62], and ResNeXt [91].
-Dataset.
-Following previous work [47], we randomly take 100 images from the NIPS17 dataset***https://www.kaggle.com/competitions/nips-2017-non-targeted-adversarial-attack.
-All images are resized into 224 $\times$ 224.
-We select 5 categories as target labels, including cat, car, ship, deer and bird.
-Hyperparameters.
-We set the perturbation budget $\epsilon=8/255$ under $\ell_{\infty}$ norm.
-We take 40 iteration steps of SSA-CWA, while other hyperparameter settings are the same as in [9].
-We set the loss function $\mathcal{L}$ to be the cross-entropy loss between the output logit $f(\bm{x})$ and the label $\bm{y}$.
-Metrics. We leverage attack success rate (ASR) as our metric, which is (adversarial examples that induce successful attacks /// all adversarial examples).
-Experiment Results.
-The results of ensemble attacks are shown in Fig.
-3.
-From our observations, we empirically derive the scaling laws between the attack success rate (ASR) and ensemble cardinality $T$ as follows:
-$\text{ASR}=\alpha\log T+C.$ (9)
-It can be explained by the following points:
-  • $C$ is influenced by various factors, such as hyperparameters, model architectures, and adversarial targets, and therefore takes different values under different conditions.
-  • For a fixed adversarial target, $\alpha$ depends solely on the model itself. Models with the same architecture share the same $\alpha$, regardless of hyperparameter settings.
-  • When $T$ is small, the measured ASR exhibits large variance due to different model selections, suggesting that our scaling laws tend to be a statistical result.
-  • ASR acts as a good representation of the empirical loss $\mathbb{L}$ defined in Eq. (2). It can be observed from Fig. 3 that there exists strong symmetry between $\mathbb{L}$ and ASR.
-  • The scaling laws become less useful as the ASR approaches 1, since ASR no longer reflects $\mathbb{L}$ at that point.
-This indicates that our observations hold only when the success rate is below the converging threshold.
-Our experiment results clearly demonstrate the universality of scaling laws, independent of specific test models or adversarial targets.
-It is important to note that both our training and evaluation models are open-sourced and pretrained on ImageNet [18], which means they do not strictly satisfy the i.i.d.
-assumption discussed in Sec.
-3.1.
-Therefore, the theorems derived in Sec.
-3.1 should be viewed as a sign of convergence and are not directly related with the empirical result measured by our experiments.
-Another key improvement observed from our scaling laws is that the evaluation loss $\mathbb{L}$ decreases exponentially with the ensemble cardinality $T$, which is much tighter compared to bounds derived in previous works [41, 94].
-3.3 Scaling Laws on Defended Models
-Figure 4: Scaling Laws for defended models. While our adversarial examples exhibit high attack success rates across defenses, clear linear scaling laws can be observed from the fitted lines.
-Figure 5: Scaling Laws for MLLMs.
-We measure the ASR of target MLLMs and fit the results with lines.
-Note that scaling laws still exist, except for the fact that results vary largely between models and targets.
-To protect neural networks from attacks, various defense mechanisms have been proposed.
-Our scaling laws work well with these defense mechanisms, showing that these laws tend to be a general result across defenses.
-Target Models.
-We select six defended models as target models, including 2 adversarially trained models – FGSMAT [52] and Ensemble AT (EnsAT) [79], and 4 input purification defenses – high-level representation guided denoiser (HGD) [56], bit depth reduction (BDR) [24], JPEG compression [16], and DiffPure [67].
-We use the same surrogate models and algorithms as Sec.
-3.2.
-Experiment Results.
-As demonstrated in Fig.
-4, we observe a clear scaling law in all these defended models, further demonstrating its universality.
-By ensembling a large number of normally trained surrogate models in Sec.
-3.2, we improve the attack success rate of DiffPure [67] by more than 30%, adversarially trained models by more than 20%, and high-level guided denoiser [56] by more than 70%.
-This demonstrates that simply increasing the number of surrogate models can break state-of-the-art defenses like DiffPure [67] and adversarial training [65].
-3.4 Scaling Laws on MLLMs
-Multimodal Large Language Models (MLLMs) [58, 68, 29] have demonstrated impressive abilities in vision understanding, captioning, and reasoning.
-In this section, we show that our scaling laws persist in these MLLMs.
-Target Models.
-We still leverage the adversarial examples generated against traditional image classifiers and select 3 popular VLMs – LLaVA-v1.5-7b [58], DeepSeek-VL-7b [17], and Phi-3-vision [29] as target models.
-Evaluation Metrics.
-As MLLMs output text rather than logits, we adopt the definition of ASR from MultiTrust [98], which leverages GPT-4 for determining whether the targeted object is stated in the description.
-In addition to assessing the correctness of the image classification, it introduces prompts to evaluate the degree of distortion in the crafted image.
-If GPT determines that the image is highly corrupted, the attack is considered unsuccessful, even if the target model misclassifies the image.
-This metric ensures that our judgments of successful attacks are of high quality.
-Experiment Results. From the results presented in Fig. 5, several key observations can be drawn:
-  • All models exhibit a similar value of $C$, indicating that they are all highly robust to attack patterns when the ensemble cardinality $T$ is small.
-  • $\alpha$ varies significantly across models, revealing the considerable differences in robustness between MLLMs.
-  • Since MLLMs are trained on multimodal data, different adversarial targets significantly influence the attack success rates.
-This variability may stem from the inherent characteristics of the target labels, as naturally some labels (such as “deer”) are less representative compared to labels like “cat”.
-This phenomenon, often referred to as class imbalance, has been discussed in prior studies [46].
-When compared to image classifiers, MLLMs demonstrate superior robustness across all chosen adversarial targets.
-This supports the general opinion that multimodal training leads to more robust representations [42], enabling the model to focus more on the essential features of the images rather than specific attack patterns.
-Furthermore, in contrast to traditional views that scaling model parameters results in increased robustness, we find that among the three target models, Phi-3-vision exhibits the highest robustness.
-This suggests that the scaling laws derived previously in [5] may not be directly applicable to multimodal models.
-4 Extending Scaling Laws to Vision Encoders
-Figure 6: Scaling Laws when we use CLIPs as surrogate models under $\epsilon=8/255$. We take ASR measured by Metric 1 to ensure our results are of high confidence.
-In this section, we extend our findings of scaling laws from image classifiers to multimodal foundation models such as CLIP [70].
-Sec.
-4.1 shows the extension of Eq.
-(2) to image embeddings, while Sec.
-4.2 demonstrates high success rates over commercial multimodal large language models (MLLMs) under our framework.
-4.1 Attacking Image Embeddings
-Contrastive Language-Image Pretraining (CLIP) is a popular self-supervised framework that can pretrain large-scale language-vision models on web-scale text-image pairs via contrastive learning [70].
-Compared to image classifiers that are normally trained on image classification datasets, CLIPs have demonstrated superior zero shot generalization abilities in multimodal downstream tasks [69, 53].
-Based on the generalization ability of CLIP models, previous works have utilized them as surrogate models in model ensembles [26, 1, 32].
-In fact, the targeted attack framework described in Eq.
-(2) can be easily extended by replacing the loss function $\mathcal{L}$ with the similarity measure function $\operatorname{sim}(\cdot)$ from CLIP (usually cosine embedding loss in practice), and by substituting the image classifiers $f$ with CLIP’s image encoders $f^{I}$.
-This modification leads to the following optimization objective:
-$\min_{\bm{x}}\frac{1}{T}\sum_{j=1}^{T}\operatorname{sim}(f^{I}_{j}(\bm{x}),f^{I}_{j}(\bm{x}_{\text{tar}})),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (10)
-Note that besides an empirical loss function, Eq.
-(10) can also be explained as minimizing the distance between the embeddings of the adversarial examples and the target images on a low-dimension embedding manifold.
-4.2 Experiments
-We first introduce our settings for attacking pretrained CLIP models and MLLMs used for our evaluation.
-Surrogate Models.
-We select 12 pretrained open-source CLIP models via OpenCLIP [12].
-In contrast to previous works that typically use fewer than 4 surrogate models, our ensemble choice is sufficient to yield robust results.
-The full list of our surrogate clips can be found in Appendix.
-E.
-Target Models.
-We select 6 commercial MLLMs as our black-box test models - GPT-4o, GPT-4 Turbo [68], Claude-3.5 Sonnet https://www.anthropic.com/news/claude-3-family, Gemini-1.5 Pro https://deepmind.google/technologies/gemini/pro/, Hunyuan [30], Qwen [28] - and 3 open-source MLLMs - LLama-3.2 90B §§§https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Pixtral Large https://mistral.ai/news/pixtral-large/, and DeepSeek-VL 7B [17].
-ASR Target: Cat Target: Car Target: Ship Target: Deer Target: Bird
-$\epsilon$ ($\ell_{\infty}$ norm) $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$ $8/255$ $16/255$
-GPT-4o 0.58 0.99 0.22 0.52 0.57 0.91 0.40 0.89 0.56 0.97
-GPT-4-Turbo 0.68 0.99 0.23 0.55 0.59 0.92 0.38 0.86 0.61 0.97
-Claude-3.5 0.56 0.91 0.18 0.42 0.41 0.79 0.37 0.74 0.38 0.80
-Gemini-1.5 Pro 0.4 0.92 0.16 0.50 0.35 0.82 0.29 0.82 0.38 0.83
-Hunyuan-Vision 0.58 0.95 0.23 0.53 0.59 0.90 0.42 0.88 0.52 0.95
-Qwen-VL Plus 0.75 0.98 0.28 0.54 0.65 0.91 0.52 0.92 0.69 0.97
-Llama-3.2 90B 0.48 0.88 0.21 0.48 0.42 0.87 0.39 0.87 0.4 0.86
-Pixtral Large 0.49 0.88 0.19 0.49 0.46 0.87 0.36 0.77 0.44 0.87
-Deepseek-VL 7B 0.54 0.96 0.17 0.52 0.52 0.87 0.44 0.91 0.51 0.93
-Table 2: ASR under Metric 2 measured across different settings of $\epsilon$.
-Remarkably, combined with our methods, most adversarial examples are considered to contain the targets’ features by MLLMs given $\epsilon=16/255$.
-Hyperparameters.
-To demonstrate the scaling laws, we set $\epsilon=8/255$.
-For practical applications, we follow the common settings in [26, 64] and set $\epsilon=16/255$.
-All other experiment settings are consistent with those in Sec.
-3.
-Evaluation Metrics.
-We employ two types of evaluation metrics in this study.
-The first metric is the same as in Sec.
-3.4.
-The second metric focuses on the features of adversarial examples, whose goal is to evaluate the insertion of features within adversarial targets into these examples.
-It requires the target model to generate a description and asks GPT whether the description mentions the adversarial target.
-Both metrics take ASR as the final score, which we refer to as Metric 1 and Metric 2 in the following sections for simplicity.
-Test prompts can be found in Appendix D.
-Experiment Results.
+We select 7 state-of-the-art MLLMs as our black-box targets, including 4 commercial models (GPT-4o [openai2024gpt4technicalreport], Claude-3.5-Sonnethttps://www.anthropic.com/news/claude-3-family, Gemini-2.5-Pro [comanici2025gemini], Doubao-Seed-1.6https://seed.bytedance.com/en/blog/introduction-to-techniques-used-in-seed1-6) and 3 state-of-the-art open-source models (Llama-3.2-90B-Vision-Instruct§§§https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/, Qwen3-VL-235B-A22B-Instructhttps://qwen.ai/blog?id=99f0335c4ad9ff6153e517418d48535ab6d8afef&from=research.latest-advancements-list, and DeepSeek-VL2 [wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels]).
+Dataset and Hyperparameters.
+We use the same SSA-CWA attack protocol as in Sec.
+3.2, but with a simplified dataset for this section.
+We randomly sample 100 images from the NIPS17 dataset with the same 8 distinct target classes, resulting in 800 unique source-target pairs per ensemble configuration.
+All experiments in this section are conducted using a unified perturbation budget of $\epsilon=16/255$ under the $\ell_{\infty}$ norm, a common and widely applied setting for practical attacks [dong2023robust, luo2024imageworth1000lies].
+Evaluation Metric.
+We employ a single, unified evaluation metric for all MLLM targets.
+As MLLMs produce generative outputs instead of classification logits, we adopt the LLM-as-Judge framework from MultiTrust [zhang2024benchmarking].
+We leverage GPT-4o to semantically evaluate the target model’s description.
+An attack is considered successful only if GPT-4 judges that (1) the model’s description mentions the adversarial target concept and (2) the description does not mention factors indicating the image is of low quality, noisy, or modified.
+The prompt of this strict, high-quality metric is provided in Appendix D.
+4.2 Generality and Power of Scaling Laws
+Our results, conducted on the latest generation of MLLMs, deliver two key findings.
+The log-linear scaling laws persist.
 Fig.
-6 presents the experiment results on the test models under Metric 1.
-These results reveal that scaling laws continue to hold when using CLIP models as surrogate models.
-This outcome seems natural as image features are still correctly learned within the vision encoders.
-Our method achieves success rates of over 30% on most adversarial targets under Metric 1, indicating that more than $1/3$ of the original images are transformed into high-quality adversarial examples with minimal noise, which demonstrates the efficiency of our method.
-We also conduct extensive experiments under Metric 2.
-From Table.
-2 we observe that the high attack success rates of our adversarial examples are generalized results, and the transferability of our adversarial examples is strong enough to fool most popular MLLMs.
-Compared to previous methods that tried to attack commercial MLLMs by visual adversarial examples [26], our methods exhibit a 40% higher success rate.
-This demonstrates the strength of our core inspiration of scaling, and highlights the practical applications of our methods.
-Meanwhile, we find the recently released open-source models Llama-3.2 90B and Pixtral Large exhibiting a better performance than other models generally, which indicates that the AI community is running towards developing more robust, safe and reliable MLLMs.
-5 Discussions
-In this section, we further demonstrate the scaling laws by using different optimizers, visualizing the adversarial perturbations, and collecting the fitting error.
-5.1 Scaling Laws are General Principles
-The idea that enlarging the model ensemble induces higher attack success rates has been proposed in previous work [94].
-Nevertheless, while prior work usually provides qualitative guesses or bounds, we exploit linear scaling laws by extensive quantitative experiments.
-While our scaling laws work well with targeted attack settings and the SSA-CWA algorithm, further experiment results reveal that our scaling laws are general principles.
+5 presents the ASR results on our 7 SOTA MLLM targets as a function of ensemble cardinality ($T$).
+Our first key finding is that the log-linear scaling laws, discovered in Sec.
+3.3, robustly generalize to this new domain.
+This outcome confirms that the principle is not an artifact of standard classifiers but a fundamental property of ensemble attacks.
+Even when attacking the complex, SOTA vision encoders of models, scaling the number of CLIP surrogates ($T$) yields a predictable, linear increase in ASR on a log scale.
+Scaling is a powerful and practical weapon.
+Having verified the scaling laws (Fig.
+5), we now examine its practical impact.
+Table 1 reports the final ASR from our full scaled ensemble ($T=12$).
+The results are twofold: they confirm the devastating power of scaling while simultaneously revealing a stark hierarchy in SOTA model robustness.
+The attack’s power is most evident on Qwen3-VL-235B and GPT-4o, which succumb to a 90.0% and 85.0% average ASR, respectively.
+This performance represents a significant leap in practical capability, surpassing the 40-50% ASRs reported by previous methods on commercial models [dong2023robust].
+However, this effectiveness is not universal.
+Claude-3.5-Sonnet exhibits exceptional robustness, neutralizing the attack to a mere 31.0% ASR.
+This resilience stands in sharp contrast to its direct competitor, Gemini-2.5-Pro, which is significantly more vulnerable at 69.0% ASR.
+The open-source landscape mirrors this variability, with DeepSeek-VL2 (59.0%) and Llama-3.2-90B (58.0%) settling into a moderate robustness tier.
+These results demonstrate that scaling is a powerful and practical tool, and the resulting ASRs serve as a clear benchmark for the diverging robustness of modern MLLMs.
+Target Model - Label Sea Lion Barometer Swing Pickelhaube Fig Lock Computer Snake Average
+   GPT-4o 0.86 0.76 0.73 0.92 0.84 0.86 0.93 0.92 0.85
+   Claude-3.5-Sonnet 0.25 0.19 0.34 0.38 0.38 0.14 0.38 0.41 0.31
+   Gemini-2.5-Pro 0.72 0.44 0.57 0.79 0.86 0.65 0.78 0.68 0.69
+   Doubao-Seed-1.6 0.61 0.73 0.76 0.87 0.78 0.70 0.69 0.62 0.72
+   Llama-3.2-90B-Vision 0.54 0.45 0.65 0.70 0.54 0.55 0.61 0.61 0.58
+   Qwen3-VL-235B 0.90 0.93 0.80 0.95 0.95 0.91 0.98 0.79 0.90
+   DeepSeek-VL2 0.68 0.35 0.63 0.50 0.63 0.67 0.54 0.68 0.59
+Table 1: Attack Success Rate (ASR) on SOTA MLLMs using our full scaled ensemble ($T=12$) and practical budget ($\epsilon=16/255$).
+The ASR is measured using our unified LLM-as-Judge metric.
+We report the per-target ASR and the average ASR across all 8 target classes.
+5 Further Analysis and Discussion
+In this section, we further demonstrate the robustness of our discovered scaling laws through ablation studies (Sec.
+5.1) and provide a deeper analysis into the semantic nature of the perturbations generated via scaling (Sec.
+5.2).
+5.1 Ablations on the Scaling Laws’ Generality
+While our main experiments (Sec.
+3) focused on targeted attacks using SSA-CWA, here we verify that the log-linear scaling laws are general principles that hold under different attack settings and across various optimizers.
 (a) Under untargeted setting.
-(b) Across different algorithms.
-Figure 7: Ablations on untargeted settings and other algorithms.
-We mainly follow the experiment configurations in Sec.
-3.1.
-For untargeted attacks, we set $\epsilon=2/255$.
-For other algorithms, we leverage cat as the target class and plot the average ASR over all target models.
-(a) It demonstrates our linear scaling laws work under untargeted settings.
-(b) It supports our claim that linear trends occur on variance-reduction algorithms, while MI-FGSM exhibits inferior performance due to gradient conflicts.
-Untargeted Attack Setting.
-Untargeted attacks are similar to targeted attacks except for a different training objective from Eq.
-(2), and all our algorithms are obviously compatible with the untargeted setting:
-$\max_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}_{\text{nat}}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (11)
-Our experiment results in Fig.
-7(a) serve as a clear sign that scaling laws remain valid over untargeted settings.
-We mainly follow our settings in Sec.
-3.1, while we set $\epsilon=2/255$, and measure the results on 3 target models.
-Taking an extremely small $\epsilon$ is necessary for demonstrating our linear scaling laws, otherwise the attack success rate will easily reach $100\%$ and therefore come to a plateau.
-The results validate our claim that scaling laws exist regardless of specific attack targets.
-Across Different Attack Algorithms.
-Besides SSA-CWA, we also provide experimental validations of different attack algorithms in Fig.
-7(b).
-As stated in Sec.
-2.2, severe gradient conflicts can be observed when we iteratively update the input image $\bm{x}$ if we directly perform backward propagation over the sum of output logits (e.g., as in MI-FGSM).
-Gradient conflicts prevent the input $\bm{x}$ from being updated properly and lead to missing information of model weaknesses.
-However, advanced algorithms, such as SVRE [92] and CWA [9] address this issue by leveraging optimizers that can reduce the variance of gradients.
-Our experiment results support our claim that scaling laws occur only on such algorithms.
-Note that the attack success rate of MI-FGSM also becomes higher with increasing ensemble cardinality, further validating that our theoretical insight from Theorem 2 is regardless of specific attack methods.
-5.2 Visualizations of Adversarial Perturbations
-Figure 8: Visualization of adversarial perturbations. A clear trend of the increasing semantic interpretability of adversarial perturbations over different target objects.
+(b) Across different optimizers.
+Figure 6: Ablations on the generality of the log-linear scaling laws.
+(a) The laws persist in the untargeted setting (Eq.
+3).
+(b) The laws hold for advanced optimizers (MI-CWA, VMI-CWA, MI-SVRE) that mitigate gradient conflict.
+In contrast, the naïve logit ensemble strategy fails to scale, supporting our hypothesis in Sec.
+$2.$
+5.1.1 Untargeted Attack Setting.
+We first validate the scaling laws in the untargeted setting (defined by Eq.
+3).
+To do this, we follow the core experimental settings from Sec.
+3.2, using the SSA-CWA algorithm on the same 64-model surrogate pool and NIPS17 dataset.
+We measure the average ASR against 3 representative target models from Family 1 (Swin-B, ConvNeXt-B, and ResNeXt50).
+As the untargeted setting is a significantly simpler task, We use a small $\epsilon=2/255$ to avoid the ASR quickly saturating at 100%, which would obscure the underlying trend.
 As shown in Fig.
-8, the semantic meaning of our adversarial perturbations also increases as the cardinality of the model ensemble grows.
-This suggests that the attack success rates are closely related to the interpretability of the perturbations, aligning with human observations and preferences [15].
-Since the ultimate goal of adversarial attacks is to generate high-quality adversarial examples, it is preferable to have interpretable perturbations, as this reinforces the idea that the scaling effectively helps learn common features rather than merely converging to local optima.
-As people traditionally believe images lie on a low dimensional manifold of desirable properties [31], this interpretability of adversarial examples can be also understood within the feature space of images, where human-designed attack algorithms aim to eliminate the non-robust features of images and shift the image towards the robust features of the target class [45].
-From this perspective, our approach of enlarging model ensembles seeks to provide more information about the manifold itself, offering more accurate representations of the image features.
-When we enlarge $\epsilon$ from $8/255$ to $16/255$ or even larger, the effect of the natural image goes down to 0, and the adversarial attack task degenerates to an image generation task.
-We claim that our method of scaling model ensemble still performs well under the degenerated image generation task, where the visualized results are shown in Appendix F.
-5.3 Fitting Error of Scaling Law Curve
-To comprehensively demonstrate the universality of our scaling laws, we provide the MSE fitting error from Fig.
-3 in Table.3.
-It can be observed that the small fitting error acts as a quantitative validation of our scaling laws.
-Thus, our scaling laws remain a scalable way to enhance the transferability of adversarial examples.
-Targets Cat Car Ship Deer Bird
-Metric ASR Loss ASR Loss ASR Loss ASR Loss ASR Loss
-Fitting Error 0.0040 0.34 0.0040 0.17 0.0038 0.46 0.0025 0.05 0.0023 0.27
-Table 3: The averaged MSE fitting error of our curves in Fig.
-3.
-Note that ASR usually has range $[0,1]$, while loss has range $[0,10]$.
-The fitting error is averaged on all test models, not only the test models that are fitted with lines for demonstration.
+6(a), even under untargeted conditions, a clear log-linear trend persists.
+This confirms that the scaling laws are a general principle of ensemble attacks, independent of the specific attack objective.
+5.1.2 Across Different Attack Algorithms.
+We now provide a tailored experiment to validate our hypothesis from Sec.
+2.2: that the log-linear scaling laws are unlocked only by advanced ensemble optimizers (like CWA or SVRE) that resolve gradient conflict, while naïve ensemble strategies fail to scale.
+Experimental Design.
+To test this, we design two families of attack algorithms and evaluate their scaling performance.
+All experiments follow the protocol in Sec.
+3.2 (targeting sea lion) and results are averaged over our standard target classifiers (Family 1 of Sec.
+3.2).
+  • Family 1 (Logit Ensemble Strategies): We test the logit averaging strategy, which we hypothesize will fail to scale.
+We combine the strategy with various powerful gradient optimizers and input transformations: MI-FGSM [dong2018boosting], DI-MI-FGSM [xie2019improving], and VMI-FGSM [wang2021enhancing].
+  • Family 2 (Advanced Ensemble Strategies): We then test advanced gradient optimizers.
+This family includes MI-CWA [chen2023rethinking], VMI-CWA [wang2021enhancing, chen2023rethinking], and MI-SVRE [xiong2022stochastic].
+Results and Analysis.
+The results are shown in Fig.
+6(b).
+All algorithms in Family 1, which use logit averaging, fail to exhibit the log-linear scaling law.
+While performance can be boosted by components like DI or VMI, their ASR stagnates at a low value.
+Conversely, algorithms in Family 2 demonstrate a clear and persistent log-linear scaling law.
+This behavior mirrors the results of our main SSA-CWA algorithm (from Sec.
+3), confirming that the scaling phenomenon is robust across different advanced optimizers, so long as they resolve the conflict.
+This experiment confirms our claim: the ability to scale an ensemble attack is contingent on the ensemble optimization strategy.
+5.2 The Semantic Nature of Scaled Perturbations
+Figure 7: Visualization of adversarial perturbations as $T$ scales.
+We take 5 labels from CIFAR10 [krizhevsky2009learning] to create adversarial examples as they are relatively easy for human eyes to recognize.
+As the ensemble cardinality grows (top to bottom), the perturbations evolve from unstructured noise ($T=1$) into semantically meaningful patterns that resemble features of target classes ($T=64$).
+Beyond attack success rates, we investigate what the ensemble attack learns as $T$ increases.
+As shown in Fig.
+7, scaling the ensemble fundamentally changes the nature of the resulting perturbation.
+At $T=1$, the perturbation is noisy and model-specific.
+However, as $T$ increases to 64, the perturbation evolves to contain clear, interpretable, and semantic features of the target class (e.g., textures for ‘Cat’, an owl-like shape for ‘Bird’).
+This is a critical finding that provides a qualitative explanation for the scaling laws.
+First, scaling $T$ forces the optimizer to discard model-specific, ‘non-robust features’ [ilyas2019adversarialexamplesbugsfeatures] and instead find a ‘common weakness’ solution that fools all models.
+Next, our visualization strongly suggests this common weakness is not just a mathematical artifact but is tied to the semantic essence of the target class.
+The attack learns to inject robust features of the target [dai2022fundamental].
+This reinforces that our scaling approach is a principled method to distill the shared, transferable features of a concept from a diverse set of models.
+As noted by [fawzi2018adversarial], images lie on a low-dimensional manifold; our scaling approach appears to provide a more accurate and robust representation of this manifold for the target class.
 6 Conclusion
-In this work, we investigate the scaling laws for black-box adversarial attacks.
-Theoretically, we provide asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec.
-3.1.
-Empirically, we show that the attack success rate and the loss are converging exponentially.
-We verify this by comprehensive experiments over traditional image classifiers (Sec.
-3.2), defended models like AT and DiffPure (Sec.
-3.3), and modern MLLMs (Sec.
-3.4 and Sec.
-4) using various optimizers (Sec.
-5).
-These experimental results demonstrate the universality of scaling laws in transfer attacks.
-Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than focusing on designing intricate algorithms on a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy multimodal large foundation models.
+In this work, we investigate scaling laws for black-box adversarial attacks.
+We find that while gradient conflict stalls naïve ensembling, advanced optimizers (e.g., CWA) unlock a robust log-linear scaling law: the Attack Success Rate increases linearly with $\log T$.
+We empirically verify this law’s universality across standard classifiers (Sec.
+3.3), SOTA defenses , and MLLMs (Sec.
+3.4), and show that scaling distills semantic features (Sec.
+5.2).
+Applying this methodology to SOTA MLLMs (Sec.
+4) serves as a powerful benchmark.
+The scaled attack achieves 85.0% ASR on GPT-4o, while Claude-3.5-Sonnet shows exceptional resilience at only 31.0% ASR.
+Our findings urge a shift in focus for robustness evaluation: from designing intricate algorithms on small ensembles to understanding the principled and powerful threat of scaling, which is crucial for building and verifying genuinely robust foundation models.
 Appendix A Proof of Theorem 1.
@@
 $\sqrt{T}\nabla\hat{\mathbb{L}}(\bm{x}^{*})$ $=\sqrt{T}(\nabla\hat{\mathbb{L}}(\bm{x}^{*})-\nabla\mathbb{L}(\bm{x}^{*}))$
-[…]
+$\overset{d}{\to}\mathcal{N}(0,\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*})))),$
 and the law of large numbers:
-[…]
+$\nabla^{2}\hat{\mathbb{L}}(\bm{x}^{*})\overset{p}{\to}\nabla^{2}\mathbb{L}(\bm{x}^{*}).$
 Plug into the previous equation using $A\mathcal{N}(0,\Sigma)=\mathcal{N}(0,A\Sigma A^{T})$:
-[…]
-[…]
+$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\overset{d}{\to}(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}\mathcal{N}(0,\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*}))))$
+$\overset{d}{=}\mathcal{N}(0,\nabla^{2}\mathbb{L}(\bm{x}^{*})^{-1}\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*}))\nabla^{2}\mathbb{L}(\bm{x}^{*})^{-1}).$
 The above indicate $\hat{\bm{x}}-\bm{x}^{*}$ converges in $\mathcal{O}(\frac{1}{\sqrt{T}})$.
@@
 We select these models based on their diversity in architectures, training methods, and training datasets.
-Results are shown in Table 4 and Table 5, where all of our models come from open-source platforms such as Torchvision [66] and Timm [84].
+Results are shown in Table 2 and Table 3, where all of our models come from open-source platforms such as Torchvision [torchvision2016] and Timm [rw2019timm].
 Note that default in the table means normally trained on ImageNet-1K dataset.
@@
 We provide the pseudo-code version of SSA-CWA implementation in Algorithm 1.
-Their combination is induced from  [63] and  [9], while implementations of other algorithms (e.g., MI-FGSM, MI-CWA) can be found in their original papers.
+Their combination is induced from  [long2022frequency] and  [chen2023rethinking], while implementations of other algorithms (e.g., MI-FGSM, MI-CWA) can be found in their corresponding papers.
 Algorithm 1 SSA-CWA Algorithm
@@
 4:  Update $\bm{x_{t}}$ by $\bm{x}^{0}_{t}=x_{t-1}$
-5:     for i = 1 todo
-6:         for j = 1 todo $\triangleright$ Spectrum transformation
+5:  for i = 1 to N do
+6:   for j = 1 to M do $\triangleright$ Spectrum transformation
 7:     Obtain transformed image $\mathcal{T}(\bm{x}^{j}_{t,i})$
@@
 18:Return: $\bm{x}_{T-1}$
-Appendix D Test Prompts as Metrics
-As VLMs and MLLMs output prompts rather than logits, we evaluate the attack efficacy by mainly focusing on the “classification” results of the test models, which is leveraging a GPT4 assistant to judge the content of the test models’ descriptions.
-We develop Metric 1 and Metric 2, focusing on different aspects of the descriptions.
+Appendix D Evaluation Metrics Details
+As MLLMs output generative text rather than classification logits, assessing attack success requires parsing natural language descriptions.
+In this work, we employ two distinct evaluation paradigms tailored to the specific experimental goals in Sec.
+3 and Sec.
+4.
+In Sec.
+3.4.2 (Family 3: MLLMs), we evaluate the scaling laws using a string-matching approach.
+Since MLLMs often produce descriptive text rather than the precise ImageNet class names, strict exact matching can lead to false negatives.
+Therefore, we adopt a relaxed keyword map.
+An attack is considered successful if the MLLM’s output contains any of the keywords associated with the target label.
+Table 4 lists the target labels used in our experiments and their corresponding relaxed keywords.
+Target Label (ImageNet) Accepted Keywords (Case-Insensitive)
+Sea Lion sea lion, seal, marine mammal
+Barometer barometer, gauge, measuring instrument, clock
+Swing swing, playground, seat
+Pickelhaube pickelhaube, helmet, spiked helmet, military hat
+Fig fig, fruit, plant
+Lock lock, padlock, dial
+Computer computer, pc, laptop, desktop, phone, device
+Snake snake, serpent, reptile, python, cobra
+Table 4: The relaxed keyword map used for evaluating MLLMs in Sec. 3. We map the specific ImageNet class label to a set of semantically equivalent or broader keywords.
+In Sec.
+4, to rigorously benchmark SOTA MLLMs, we adopt the LLM-as-Judge framework inspired by MultiTrust [zhang2024benchmarking].
+We use GPT-4o as the judge to evaluate the quality and success of the attack.
+The evaluation is strict: an attack is successful only if the description mentions the target and does not mention artifacts of the attack (noise, distortion).
 Fig.
-9(a) and Fig.
-9(b) display our prompts for Metric 1 and Metric 2 respectively.
-Metric 1 follows the settings in previous works [98], while Metric 2 is our customized metric to capture image features, as stated in Sec.
-4.2.
-By comparing the results in Fig.
-6 and Table.
-2, Metric 2 have a 10% higher success rate than Metric 1 under $\epsilon=8/255$ cases.
-This phenomenon is natural since some images of high noise will be still considered as success in Metric 2, which is more aligned with our practical approaches [26].
-(a) Prompts for Metric 1.
-(b) Prompt for Metric 2.
-Figure 9: Prompts of two metrics.
-(a) Besides prompt to test the adversarial example contain a target label, we also use a prompt to judge the image quality, ensuring our successful attacks are all of small noise and highly transferable.
-(b) The prompt concentrates on semantic meanings of the image descriptions.
-To consider the attack successful, it only needs the target to appear in the description.
+8 presents the exact system prompt used in our experiments.
+Prompt for GPT-4o Judge Correctness: Given a chatbot’s description of a scene: [response].
+If it is stated in the description that there is a/an [label]?
+Please answer ’Yes.’ or ’No.’.
+Prompt for GPT-4o Judge Image Quality: Given a chatbot’s description of an image: [response].
+Does the description mention factors that affect the judgment of the image content, such as low image quality, noise, or modification?
+Please answer ’Yes.’ or ’No.’.
+Figure 8: The GPT-4o system prompt used for the unified evaluation metric in Sec.
+4.
+This metric ensures that successful attacks are both effective (semantics changed) and stealthy (no obvious noise detected by the model).
 Appendix E Experiment Settings for Attacking MLLMs
@@
 Table 5: List of CLIPs that are selected as model ensemble.
-Figure 10: Visualizations of adversarial examples under $\epsilon=64/255$.
-The adversarial examples capture basic information of the target class, such as the fur of cats, tires of vehicles, cargo on ships, antlers of deer, and feathers of birds.
-In Table. 6 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [70], OpenCLIP [12], and Timm [84].
-Appendix F Degenerated Image Generation Task
-As stated in Sec.
-5.2, when $\epsilon$ goes larger, the adversarial attack task degenerates to an image generation task.
-Previous works [9, 1, 82] do not ever think about this field, but we believe that adversarial attack task and image generation task are indeed homogenous.
-We display our adversarial examples generated from 64 image classifiers (stated in Appendix B) under $\epsilon=64/255$ in Fig.
-10.
-From the visualizations we observe that our methods can capture the key features of the target class and reproduce them based on the original textures of images.
-This visualization further supports our methods are really attacking common weakness instead of running into local optima.
-We leave further explorations and analyses as future work.
-Appendix G Related Work
-G.1 Black-Box Attacks
-Deep learning models are widely recognized for their vulnerability to adversarial attacks [73, 35].
-Extensive research has documented this susceptibility, particularly in image classification contexts [2, 7], under two principal settings: white-box and black-box attacks.
-In the white-box setting, the adversary has complete access to the target model [65].
+In Table.
+5 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [radford2021learningtransferablevisualmodels], OpenCLIP [cherti2023reproducible], and Timm [rw2019timm].
+Appendix F Related Work
+F.1 Black-Box Attacks
+Deep learning models are widely recognized for their vulnerability to adversarial attacks [szegedy2013intriguing, goodfellow2014explaining].
+Extensive research has documented this susceptibility, particularly in image classification contexts [athalye2018obfuscated, carlini2017towards], under two principal settings: white-box and black-box attacks.
+In the white-box setting, the adversary has complete access to the target model [madry2017towards].
 Conversely, in the black-box setting, such information is obscured from the attacker.
-With limited access to the target model, black-box attacks typically either construct queries and rely on feedback from the model [25, 44] or leverage the transferability of predefined surrogate models [60].
+With limited access to the target model, black-box attacks typically either construct queries and rely on feedback from the model [dong2021query, ilyas2018black] or leverage the transferability of predefined surrogate models [liu2016delving].
 Query-based attacks often suffer from high time complexity and substantial computational costs, making transfer-based methods more practical and cost-effective for generating adversarial examples, as they do not necessitate direct interaction with target models during the training phase.
 Within the realm of transfer-based attacks, existing methods can generally be categorized into three types: input transformation, gradient-based optimization, and ensemble attacks.
-Input transformation methods employ data augmentation techniques before feeding inputs into models, such as resizing and padding [89], translations [23], or transforming them into frequency domain [63].
-Gradient-based methods concentrate on designing optimization algorithms to improve transferability, incorporating strategies such as momentum optimization [22, 57], scheduled step sizes [33], and gradient variance reduction [81].
-G.2 Ensemble Attacks
-Inspired by model ensemble techniques in machine learning [87, 19, 34], researchers have developed sophisticated adversarial attack strategies utilizing sets of surrogate models.
-In forming an ensemble paradigm, practitioners often average over losses [22], predictions [60], or logits [22].
-Additionally, advanced ensemble algorithms have been proposed to enhance adversarial transferability [55, 9, 92, 8].
+Input transformation methods employ data augmentation techniques before feeding inputs into models, such as resizing and padding [xie2019improving], translations [dong2019evading], or transforming them into frequency domain [long2022frequency].
+Gradient-based methods concentrate on designing optimization algorithms to improve transferability, incorporating strategies such as momentum optimization [dong2018boosting, lin2019nesterov], scheduled step sizes [gao2020patch], and gradient variance reduction [wang2021enhancing].
+F.2 Ensemble Attacks
+Inspired by model ensemble techniques in machine learning [wortsman2022model, dietterich2000ensemble, gontijo2021no], researchers have developed sophisticated adversarial attack strategies utilizing sets of surrogate models.
+In forming an ensemble paradigm, practitioners often average over losses [dong2018boosting], predictions [liu2016delving], or logits [dong2018boosting].
+Additionally, advanced ensemble algorithms have been proposed to enhance adversarial transferability [li2023making, chen2023rethinking, xiong2022stochastic, chen2023adaptive].
 However, these approaches tend to treat model ensemble as a mere technique for improving transferability without delving into the underlying principles.
-Prior studies [41] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization.
-Recent research [94] has theoretically defined transferability error and sought to minimize it through model ensemble.
-Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [59].
-G.3 Scaling Laws
+Prior studies [huang2023t] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization.
+Recent research [yao2025understandingmodelensembletransferable] has theoretically defined transferability error and sought to minimize it through model ensemble.
+Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [liu2019deep].
+F.3 Scaling Laws
 In recent years, the concept of scaling laws has gained significant attention.
-These laws describe the relationship between model size – measured in terms of parameters, data, or training FLOPs – and generalization ability, typically reflected in evaluation loss [48].
-Subsequent studies have expanded on scaling laws from various perspectives, including image classification [3, 38], language modeling [71] and mixture of experts (MoE) models [30], highlighting the fundamental nature of this empirical relationship.
-In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [50, 5].
+These laws describe the relationship between model size – measured in terms of parameters, data, or training FLOPs – and generalization ability, typically reflected in evaluation loss [kaplan2020scalinglawsneurallanguage].
+Subsequent studies have expanded on scaling laws from various perspectives, including image classification [bahri2024explaining, hestness2017deep], language modeling [sharma2022scaling] and mixture of experts (MoE) models [sun2024hunyuanlargeopensourcemoemodel], highlighting the fundamental nature of this empirical relationship.
+In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [krizhevsky2009learning, bartoldson2024adversarialrobustnesslimitsscalinglaw].
 However, the broader concept of scaling remains underexplored.
-While some preliminary observations on the effect of enlarging model ensembles can be found in prior studies [94], these effects have not been systematically analyzed.
-Recent work Wei et al.
-[83] also explore the scaling law of attack success rate when jailbreaking large language models [99, 10] by increasing number of in-context demonstrations In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models.
+While some preliminary observations on the effect of enlarging model ensembles can be found in prior studies [yao2025understandingmodelensembletransferable], these effects have not been systematically analyzed.
+Recent work wei2023jailbreak also explore the scaling law of attack success rate when jailbreaking large language models [zou2023universal, chen2025towards] by increasing number of in-context demonstrations In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models.
 Motivated by this perspective, we unveil vulnerabilities within commercial LLMs by targeting their vision encoders.